### v1

# On decoding extended Han-Zhang codes

Abstract
Extended Han-Zhang codes are a class of linear codes where each code is either a non-generalized Reed-Solomon (non-GRS) maximum distance separable (MDS) code or a near MDS (NMDS) code. They have important applications in communication, cryptography, and storage systems. While many algebraic properties and explicit constructions of extended Han-Zhang codes have been well studied in the literature, their decoding has been unexplored. In this paper, we focus on their decoding problems in terms of $\ell$-error-correcting pairs ($\ell$-ECPs) and deep holes. On the one hand, we determine the existence and specific forms of their $\ell$-ECPs, and further present an explicit decoding algorithm for extended Han-Zhang codes based on these $\ell$-ECPs, which can correct up to $\ell$ errors in polynomial time, with $\ell$ about half of the minimum distance. On the other hand, we determine the covering radius of extended Han-Zhang codes and characterize two classes of their deep holes, which are closely related to the maximum-likelihood decoding method. By employing these deep holes, we also construct more non-GRS MDS codes with larger lengths and dimensions, and discuss the monomial equivalence between them and the well-known Roth-Lempel codes. Some concrete examples are also given to support these results.
Index Terms:
Decoding, error-correcting pair, deep hole, non-GRS MDS code, NMDS code
I Introduction
Maximum distance separable (MDS) and near MDS codes are two special types of codes in coding theory, which have recently garnered significant attention due to their elegant algebraic structures and wide-ranging applications in distributed storage systems [6], random error channels [40], informed source and index coding problems [47], and secret sharing schemes [41, 55]. Formally, an $n,k,d]_{q}$ linear code $\mathcal{C}$} is a $k$-dimensional linear subspace of $\mathbb{F}_{q}^{n}$ with minimum distance $d$, and it satisfies the Singleton bound: $d\leq n-k+1$ [21], where $\mathbb{F}_{q}$ is the finite field of size $q$, $\mathbb{F}_{q}^{n}$ is the $n$-dimensional vector space, and $q=p^{m}$ is a prime power. A code $\mathcal{C}$} achieving this Singleton bound, $i.e.$, d=nk+1d=n-k+1, is called a maximum distance separable (MDS) code. If d=nkd=n-k, then $\mathcal{C}$} is referred to as an almost MDS (AMDS) code. Moreover, a code $\mathcal{C}$} is called a near MDS (NMDS) code if both $\mathcal{C}$} and $\mathcal{C}$$^{\perp}$ are AMDS codes, where $\mathcal{C}$$^{\perp}$ is the dual code of $\mathcal{C}$} with respect to a certain inner product.
I-A Linear codes where each code is either non-GRS MDS or NMDS and their decoding problems
Generalized Reed-Solomon (GRS) codes are a special class of MDS codes that have garnered a lot of attention due to their highly efficient encoding and decoding algorithms. Moreover, $n,k,n-k+1]_{q}$ GRS codes satisfy $n\leq q$, and together with the MDS conjecture $n\leq q+2$ [21], they realize nearly all possible parameter sets of MDS codes, except for the case $n>q$. However, it is still very interesting and vital to construct MDS codes that are not monomially equivalent to GRS codes, referred to as non-GRS MDS codes. These non-GRS MDS codes are important in classifying MDS codes, and are known to exhibit significant potential in resisting cryptographic attacks such as the Sidelnikov–Shestakov and Wieschebrink attacks, unlike GRS codes [27, 2]. Consequently, it is of particular interest to construct new families of linear codes in which each code is either a non-GRS MDS code or an NMDS code.
There has been a substantial body of work [32, 33, 48, 17, 19, 13, 20, 57, 14, 43, 18, 45, 46, 31, 7, 54, 23, 28, 49] focusing on the construction of new families of codes that include both MDS and NMDS codes. In particular, the (+)-twisted generalized Reed-Solomon (TGRS) codes in [19], the ($\ast$)-TGRS codes in [20], the (+)-extended TGRS ((+)-ETGRS) codes in [57], and the extended Han–Zhang codes in [31] coincide with examples of linear code families where each code is either a non-GRS MDS code or an NMDS code.
Of the codes in the previous paragraph, except for the extended Han–Zhang codes, decoding algorithms have been studied for the other three families of linear codes. Specifically, Beelen etal.et~al. [3] initially proposed a brute-force decoding method based on the established decoding algorithm of GRS codes when introducing TGRS codes. A simpler version was later presented in [4]. Although these decoding approaches are theoretically applicable for decoding (+)-TGRS and ($\ast$)-TGRS codes, they are not practical in real-world scenarios. As stated in [3] and [4], the former suffers from unexpected decoding complexity, while the latter lacks a rigorous proof ensuring that decoding works for any error vector up to the maximal decoding radius. Alternatively, He etal.et~al. [16] explored the existence of error-correcting pairs (ECPs) for (+)-TGRS and ($\ast$)-TGRS codes, and developed a corresponding decoding algorithm that operates in polynomial time. It is well-known that deep holes of linear codes are closely related to their maximum-likelihood decoding algorithms. Leveraging this insight, Fang etal.et~al. further identified the standard deep holes of (+)-TGRS codes of arbitrary lengths and explored the problem of determining all deep holes for full-length (+)-TGRS codes in [12]. Recently, Li etal.et~al. theoretically studied both the existence of ECPs and deep holes for (+)-ETGRS codes in [33]. Additionally, Jia etal.et~al. [22], Sui etal.et~al. [42], and Sun etal.et~al. [44] applied the Berlekamp–Massey algorithm for GRS codes and the extended Euclidean algorithm to propose efficient decoding algorithms for the dual codes of (+)-TGRS and ($\ast$)-TGRS codes. It is also worth noting from [19] and [20] that the dual codes of (+)-TGRS and ($\ast$)-TGRS codes may not be (+)-TGRS or ($\ast$)-TGRS codes again.
I-B Our motivations for decoding extended Han-Zhang codes based on ECPs and deep holes
From the above discussions, we can immediately summarize the following three important facts:
  • Fact 1: The current known families of linear codes where each code is either a non-GRS MDS code or an NMDS code mainly stem from TGRS codes and their extended codes.
  • Fact 2: For the aforementioned Fact 1 codes, decoding algorithms based on ECPs and deep holes appear to be more practical and versatile than other known approaches.
  • Fact 3: Although the algebraic properties of extended Han–Zhang codes have been extensively studied in [31], their decoding problems remain unexplored. In particular, it was also proved in [31] that an extended Han-Zhang code is generally not monomially equivalent to an (+)-ETGRS code, due to different weight distributions.
The ECP decoding algorithm was introduced by Kötter [26] and Pellikaan [36] in 1992 (see Definition 5). It has also been demonstrated that ECPs exist for a wide range of linear codes, including cyclic codes, GRS codes, TGRS codes with single or double twists, lengthened GRS codes, Goppa codes, alternant codes, and algebraic geometry codes [10, 26, 36, 37, 15, 16, 17, 50, 33]. In particular, if a linear code possesses an $\ell$-ECP, this framework enables a universal algebraic decoding capable of correcting up to $\ell$ errors in polynomial time.
For another decoding algorithm related to deep holes (see Definition 6), it indeed represents an extreme case of the important maximum-likelihood decoding method and is consequently both interesting and difficult. By analyzing the structure of deep hole cosets, parameters of extended codes, connections with problems and results in finite geometry, syndromes, subset sum problems, full automorphism groups, and character sums, researchers explored deep holes of first-order Reed-Muller codes, GRS codes, (+)-TGRS codes, elliptic curve codes, and dual codes of Roth-Lempel codes in [35, 53, 9, 24, 25, 29, 58, 12, 33, 48] and the references therein. Furthermore, Pellikaan etal.et~al. explored the application of ECPs in the McEliece public-key cryptosystem for quantum-resistant cryptography [38], while Yang etal.et~al. [52] and Li etal.et~al. [33] respectively leveraged deep holes of binary BCH codes and ETGRS codes to derive new binary optimal codes and non-GRS MDS codes with larger lengths and dimensions. In consideration of the above facts and applications, an important and intriguing problem naturally arises:
Problem 1.
How can we decode extended Han-Zhang codes, particularly by leveraging their ECPs and deep holes? More specifically, is it possible to develop an efficient decoding algorithm for extended Han-Zhang codes based on their ECPs, and can we identify some of their deep holes?
I-C Our contributions
We answer the aforementioned Problem 1 affirmatively, and our main contributions can be summarized as follows.
  • 1)
    We use ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ to denote an $n+1,k]_{q}$ extended Han-Zhang code (see Definition 4 for more details). Let $\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor$, where $d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))$ denotes the minimum distance of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. Utilizing structural properties, we identify extended Han-Zhang codes as subcodes of certain GRS codes and EGRS codes in Lemma 12. Building on this, we prove in Theorem 13.1) that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has only $(\ell-1)$-ECPs if it is an $n+1,k,n-k+2]_{q}$ MDS code and $2\nmid(n-k)$. We also show that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has $\ell$-ECPs in other cases by giving explicit forms of these ECPs directly in Theorems 13.2) and 14. More details on the exact values of $\ell$ and explicit forms of ECPs are summarized in Table I. Furthermore, we provide a decoding algorithm for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ based on their ECPs in Theorem 15 and Algorithm 1, supported by specific decoding examples shown in Examples 1 and 2.
  • 2)
    We prove that the covering radius of any extended Han-Zhang code is nk+1n-k+1 in Theorem 16 and establish a necessary and sufficient condition for a received vector to be a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ in Theorem 17. Leveraging these results, we further characterize when two classes of received vectors become deep holes of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ in Theorems 19 and 20, where the first class is associated with the polynomial $g(x)=g_{k-1}x^{k-1}+f(x)$ satisfying $g_{k-1}\in\mathbb{F}_{q}^{*}$ and $f(x)\in\mathcal{V}_{k}$, the second one is related to the polynomial $g(x)=g_{k+1}x^{k+1}+g_{k-1}x^{k-1}+f(x)$ with $g_{k+1}\in\mathbb{F}_{q}^{*}$, $g_{k-1}\in\mathbb{F}_{q}$, and $f(x)\in\mathcal{V}_{k}$, and $\mathcal{V}_{k}=\{f(x)=\sum_{i=0}^{k-2}f_{i}x^{i}+f_{k}x^{k}:~f_{i}\in\mathbb{F}_{q},~\forall~i\in\{0,\ldots,k-2,k\}\}$.
  • 3)
    Applying the deep holes obtained and the non-GRS MDS properties of extended Han-Zhang codes, we construct more $n+2,k+1,n-k+2]_{q}$ non-GRS MDS codes and analyze their monomial equivalence with the well-known Roth-Lempel codes in Theorems 21 and 22 as well as Remark 3. Notably, our results also confirm that there is a special family of Roth-Lempel codes where each code is either non-GRS MDS or NMDS, as stated in Remark 4. Additionally, we provide an efficient algorithm to generate these non-GRS MDS codes with larger lengths and dimensions in Theorem 23 and Algorithm 2, and validate our findings with specific examples in Examples 3 and 4.
This paper is organized as follows. After the introduction, Section II reviews fundamental notation and results on GRS codes, Roth-Lempel codes, extended Han-Zhang codes, ECPs, and deep holes. Section III establishes the existence and explicit forms of $\ell$-ECPs for extended Han-Zhang codes, and presents a decoding algorithm based on their $\ell$-ECPs. Section IV explores the deep holes of extended Han-Zhang codes and leverages these findings to construct more non-GRS MDS codes with larger lengths and dimensions. We also discuss the monomial equivalence between these newly derived non-GRS MDS codes and Roth-Lempel codes. Finally, Section V concludes this paper.
II Preliminaries
In this section, we recall some basic notation and results on GRS codes, Roth-Lempel codes, extended Han-Zhang codes, ECPs, and deep holes.
II-A Basic notation
From now on, we fix the following notation, unless stated otherwise.
  • $\mathbb{F}_{q}$ is the finite field of size $q$ and $\mathbb{F}_{q}^{n}$ is the $n$-dimensional vector space, where $q=p^{m}$ is a prime power.
  • For an $n,k,d]_{q}$ linear code $\mathcal{C}$}, we use $\dim({\mathcal{C}})$ and $d({\mathcal{C}})$ to denote its dimension and minimum distance, respectively.
  • For distinct elements $a_{1},\ldots,a_{n}\in\mathbb{F}_{q}$, we let ${\mathcal{S}}=\{a_{1},\ldots,a_{n}\}$ and let ${\bf u}=(u_{1},\ldots,u_{n})\in(\mathbb{F}_{q}^{*})^{n}$ with
    $u_{i}=\prod_{1\leq j\leq n,~j\neq i}(a_{i}-a_{j})^{-1}~{\rm for}~1\leq i\leq n.$
  • $\bf 0$} (resp. $\bf 1$}) is an appropriate row or column vector of all zeros (resp. ones), $\infty_{k+1}=(0,\ldots,0,1)^{T}$ is a column vector of length $k+1$, and $\mathbf{v}=(v_{1},\ldots,v_{n})\in(\mathbb{F}_{q}^{*})^{n}$ is a generic representation of its element.
  • For any vector ${\bf x}=(x_{1},\ldots,x_{n})\in\mathbb{F}_{q}^{n}$, ${\mathrm{supp}}(x)=\{i:~x_{i}\neq 0\}$ and $z(x)=\{i:~x_{i}=0\}$.
  • For any vectors ${\bf x}=(x_{1},\ldots,x_{n})\in\mathbb{F}_{q}^{n}$ and ${\bf y}=(y_{1},\ldots,y_{n})\in\mathbb{F}_{q}^{n}$, their standard product and Schur product are respectively defined by
    $\langle{\bf x},{\bf y}\rangle=\sum_{i=1}^{n}x_{i}y_{i}~{\rm and}~{\bf x}\star{\bf y}=(x_{1}y_{1},\ldots,x_{n}y_{n}).$
  • For any two linear codes $\mathcal{C}$} and $\mathcal{D}$} with the same length $n$, the dual code of $\mathcal{C}$} is given by
    ${\mathcal{C}}^{\perp}=\{\mathbf{c}^{\perp}\in\mathbb{F}_{q}^{n}:~\langle\mathbf{c},\mathbf{c}^{\perp}\rangle=0,~\forall~\mathbf{c}\in{\mathcal{C}}\}$
    and the Schur product code of $\mathcal{C}$} and $\mathcal{D}$} is defined by
    $\mathcal{C}\star\mathcal{D}=\{(c_{1}d_{1},\ldots,c_{n}d_{n}):~(c_{1},\ldots,c_{n})\in\mathcal{C}~{\rm and}~(d_{1},\ldots,d_{n})\in\mathcal{D}\}.$
  • ${\mathcal{C}}A=\{{\bf c}A:~{\bf c}\in{\mathcal{C}}\}$, where $\mathcal{C}$} is an $n,k]_{q}$ linear code and $A$ is an $n\times n$ matrix over $\mathbb{F}_{q}$.
  • ${\mathrm{diag}}(a_{1},\ldots,a_{n})$ denotes an $n\times n$ diagonal matrix with diagonal elements $a_{1},\ldots,a_{n}$ over $\mathbb{F}_{q}$.
  • $\mathbb{F}_{q}[x]_{k}=\{f(x)=\sum_{i=0}^{k-1}f_{i}x^{i}:~f_{i}\in\mathbb{F}_{q},~\forall~i\in\{0,\ldots,k-1\}\}$.
  • $\mathcal{V}_{k}=\{f(x)=\sum_{i=0}^{k-2}f_{i}x^{i}+f_{k}x^{k}:~f_{i}\in\mathbb{F}_{q},~\forall~i\in\{0,\ldots,k-2,k\}\}$.
II-B GRS codes, Roth-Lempel codes, and extended Han-Zhang codes
As mentioned before, an important class of MDS codes is the so-called generalized Reed-Solomon codes [21]. With the above notation, we have the following definition.
Definition 1.
( [21]) Let $k$ and $n$ be two positive integers satisfying $1\leq k\leq n\leq q$. A generalized Reed-Solomon (GRS) code, denoted by ${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v})$, is an $n,k,n-k+1]_{q}$ linear code defined by
${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v})=\{(v_{1}f(a_{1}),\ldots,v_{n}f(a_{n})):~f(x)\in{\mathbb{F}_{q}[x]_{k}}\}.$
By adding a coordinate to each codeword of a GRS code, we immediately get an $n+1,k,n-k+2]_{q}$ MDS code, namely, an extended GRS (EGRS) code
${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v},\infty)=\{(v_{1}f(a_{1}),\ldots,v_{n}f(a_{n}),f_{k-1}):~f(x)\in{\mathbb{F}_{q}[x]_{k}}\},$
where $f_{k-1}$ is the coefficient of $x^{k-1}$ in $f(x)$.
In general, we call $\mathcal{S}$} the evaluation-point sequence of ${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v})$ and ${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v},\infty)$. From [21], we can write generator matrices of ${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v})$ and ${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v},\infty)$ as
$G_{{\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v})}=\begin{pmatrix}v_{1}&\ldots&v_{n}\\ v_{1}a_{1}&\ldots&v_{n}a_{n}\\ \vdots&\ddots&\vdots\\ v_{1}a_{1}^{k-1}&\ldots&v_{n}a_{n}^{k-1}\\ \end{pmatrix}~{\rm and}~G_{{\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v},\infty)}=\begin{pmatrix}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-1}&\ldots&v_{n}a_{n}^{k-1}&1\\ \end{pmatrix}$, (1)
respectively.
Lemma 1.
( [11, Lemma 5] and [21, Theorem 5.3.3]) The following statements hold.
  1. 1)
    ${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v})^{\perp}={\mathrm{GRS}}_{n-k}({\mathcal{S}},\mathbf{v}^{-1}\star{{\mathbf{u}}})$.
  2. 2)
    ${\mathrm{GRS}}_{k}({\mathcal{S}},\mathbf{v},\infty)^{\perp}={\mathrm{GRS}}_{n-k+1}({\mathcal{S}},\mathbf{v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)$.
Definition 2.
( [39]) Let $k$ and $n$ be two positive integers satisfying $3\leq k\leq n\leq q$. For any $\delta\in\mathbb{F}_{q}$, a Roth-Lempel code, denoted by ${\mathrm{RL}}_{k,\delta}({\mathcal{S}})$, is an $n+2,k]_{q}$ linear code defined by
$\begin{split}{\mathrm{RL}}_{k,\delta}({\mathcal{S}})=\{(f(a_{1}),\ldots,f(a_{n}),f_{k-1},f_{k-2}+\delta f_{k-1}):~f(x)\in\mathbb{F}_{q}[x]_{k}\},\end{split}$ (2)
where $f_{k-1}$ and $f_{k}$ are the coefficients of $x^{k-1}$ and $x^{k}$ in $f(x)$, respectively. Moreover, it has a generator matrix of the form
$G_{{\mathrm{RL}}_{k,\delta}({\mathcal{S}})}=\left(\begin{array}[]{cccccc}1&\ldots&1&0&0\\ a_{1}&\ldots&a_{n}&0&0\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ a^{k-3}_{1}&\ldots&a^{k-3}_{n}&0&0\\ a^{k-2}_{1}&\ldots&a^{k-2}_{n}&0&1\\ a^{k-1}_{1}&\ldots&a^{k-1}_{n}&1&\delta\\ \end{array}\right).$ (9)
Definition 3.
For a subset ${\mathcal{S}}\subseteq\mathbb{F}_{q}$, we call $\mathcal{S}$} an $(n,k,\delta)$-set if
$|{\mathcal{S}}|=n~{\rm and}~a_{i_{1}}+\ldots+a_{i_{k}}\neq\delta~{\rm for~any}~\{a_{i_{1}},\ldots,a_{i_{k}}\}\subseteq{\mathcal{S}}.$
Moreover, $\mathcal{S}$} is an $(n,k,0)$-set if and only if $\mathcal{S}$} is $k$-zero-sum free. Conversely, we say that $\mathcal{S}$} contains a $k$-zero-sum subset if $\mathcal{S}$} is not $k$-zero-sum free.
In [39], Roth and Lempel established a necessary and sufficient condition for a Roth-Lempel code ${\mathrm{RL}}_{k,\delta}({\mathcal{S}})$ to be non-GRS MDS as follows.
Lemma 2.
( [39, Page 655]) For any $3\leq k\leq n\leq q$ and $\delta\in\mathbb{F}_{q}$, a Roth-Lempel code ${\mathrm{RL}}_{k,\delta}({\mathcal{S}})$ is an $n+2,k,n-k+3]_{q}$ non-GRS MDS code if and only if $\mathcal{S}$} is an $(n,k-1,\delta)$-set.
Definition 4.
( [31, Definition 1]) Let $n$ and $k$ be two positive integers satisfying $3\leq k\leq n-2\leq q-2$. An extended Han-Zhang code, denoted by ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,k]_{q}$ linear code
${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)=\{(v_{1}f(a_{1}),\ldots,v_{n}f(a_{n}),f_{k}):~f(x)\in\mathcal{V}_{k}\},$
where $f_{k}$ is the coefficient of $x^{k}$ in $f(x)$.
The following lemmas recall some basic properties of extended Han-Zhang codes obtained in [31].
Lemma 3.
( [31, Theorem \@slowromancapiii@.2]) The extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has a generator matrix and a parity-check matrix of the form
$G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}=\begin{pmatrix}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1\\ \end{pmatrix}~{\rm and}~H_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}=\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{n-k-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{n-k-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{n-k-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{n-k-1}&-1\\ \frac{u_{1}}{v_{1}}a_{1}^{n-k}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{n-k}&-\sum_{i=1}^{n}a_{i}\end{pmatrix}$, (10)
respectively.
Lemma 4.
( [31, Theorem \@slowromancapiii@.4]) Any extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is either MDS or NMDS. Furthermore, the following statements hold.
  1. 1)
    ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,k,n-k+2]_{q}$ MDS code if and only if $\mathcal{S}$} is $k$-zero-sum free.
  2. 2)
    ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,k,n-k+1]_{q}$ NMDS code if and only if $\mathcal{S}$} contains a $k$-zero-sum subset.
Lemma 5.
( [31, Theorem \@slowromancapiii@.7]) Any extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is non-GRS.
II-C ECPs and deep holes
This subsection reviews some basic definitions of ECPs and deep holes. For any linear code $\mathcal{C}$}, a pair of linear codes $(\mathcal{A},\mathcal{B})$ is an $\ell$-error-correcting pair ($\ell$-ECP) of $\mathcal{C}$} that can yield a decoding algorithm capable of uniquely correcting up to $\ell$ errors in polynomial time. From [36, Corollary 2.15], we have $\ell\leq\lfloor\frac{d({\mathcal{C}})-1}{2}\rfloor$. The following is an alternative definition of the $\ell$-ECP.
Definition 5.
( [10, 36]) Let $\mathcal{C}$} be an $n,k]_{q}$ linear code. Let $\mathcal{A}$ and $\mathcal{B}$ be two linear codes with the same length $n$ over $\mathbb{F}_{q}$. We call $(\mathcal{A},\mathcal{B})$ an $\ell$-error-locating pair ($\ell$-ELP) of $\mathcal{C}$} if all of following three conditions hold:
${\rm i)}~\mathcal{A}\star\mathcal{B}\subseteq{\mathcal{C}}^{\perp};~~{\rm ii)}~d(\mathcal{B}^{\perp})>\ell;~~{\rm iii)}~\dim(\mathcal{A})>\ell.$
Moreover, if an $\ell$-ELP $(\mathcal{A},\mathcal{B})$ satisfies the additional condition
${\rm iv)}~d(\mathcal{A})+d({\mathcal{C}})>n,$
we call it an $\ell$-error-correcting pair ($\ell$-ECP) of $\mathcal{C}$}.
The following three lemmas are useful.
Lemma 6.
( [17, Remark 2.1]) Let ${\mathcal{C}}_{i}$ be an $n,k_{i}]_{q}$ linear code for $i=1,2$ with ${\mathcal{C}}_{1}\subseteq{\mathcal{C}}_{2}$ and $(\mathcal{A},\mathcal{B})$ be an $\ell$-ECP of ${\mathcal{C}}_{2}$. Then the following statements hold.
  • 1)
    For any vector ${\bf x}\in(\mathbb{F}_{q}^{*})^{n}$, $({\bf x}^{-1}\star\mathcal{A},{\bf x}\star\mathcal{B})$ is an $\ell$-ECP of ${\mathcal{C}}_{2}$.
  • 2)
    $(\mathcal{A},\mathcal{B})$ is an $\ell$-ECP of ${\mathcal{C}}_{1}$.
Lemma 7.
( [34, Example 4.2] and [34, Theorem 6.2]) Let $\mathcal{C}$} be an $n,n-2\ell,2\ell+1]_{q}$ MDS code, where $\ell$ is a positive integer with $1\leq\ell\leq\frac{n}{2}-1$. Then $\mathcal{C}$} has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ if and only if $\mathcal{C}$} is a GRS code. Moreover, both $\mathcal{A}$ and $\mathcal{B}$ are GRS codes with the same evaluation-point sequence as that of $\mathcal{C}$}, and $\mathcal{B}=(\mathcal{A}\star\mathcal{C})^{\perp}$.
Lemma 8.
( [37, Proposition 2.5]) Let $\mathcal{C}$} be an $n,n-2\ell,2\ell+1]_{q}$ MDS code and let $(\mathcal{A},\mathcal{B})$ be an $\ell$-ECP of $\mathcal{C}$}. Then $\mathcal{A}$ is an $n,\ell+1,n-\ell]_{q}$ MDS code.
Definition 6.
( [21]) Let $\mathcal{C}$} be an $n,k]_{q}$ linear code. For any vector $\mathbf{u}\in\mathbb{F}_{q}^{n}$, its error distance to $\mathcal{C}$} is defined as $d(\mathbf{u},{\mathcal{C}})=\min\{d(\mathbf{u},\mathbf{c}):~\mathbf{c}\in{\mathcal{C}}\}$. We call $\rho({\mathcal{C}})=\max\{d(\mathbf{u},{\mathcal{C}}):~\mathbf{u}\in\mathbb{F}_{q}^{n}\}$ the covering radius of $\mathcal{C}$} and the deep holes of $\mathcal{C}$} are just the vectors in $\mathbb{F}_{q}^{n}$ that achieve $\rho({\mathcal{C}})$.
The following lemma provides a way to get more non-GRS MDS codes from a known non-GRS MDS code with respect to some conditions.
Lemma 9.
( [33, Lemma 2.4]) Let $\mathcal{C}$} be an $n+1,k,n-k+2]_{q}$ MDS code with generator matrix $G$ and covering radius $\rho({\mathcal{C}})=n-k+1$. Then for any $\mathbf{u}\in\mathbb{F}_{q}^{n+1}$, $\mathbf{u}$ is a deep hole of $\mathcal{C}$} if and only if the matrix $\left(\left(\begin{array}[]{c}G\\ \mathbf{u}\end{array}\right),~\infty_{k+1}\right)$ generates an $n+2,k+1,n-k+2]_{q}$ MDS code $\mathcal{C}$$^{\prime}$. Moreover, if $\mathcal{C}$} is non-GRS, so is $\mathcal{C}$$^{\prime}$.
The following two lemmas are also relevant in identifying deep holes of extended Han-Zhang codes.
Lemma 10.
([1, Lemma 7.3]) Let $\mathcal{C}$} be an $n,k]_{q}$ linear code with generator matrix $G$. Then $\mathcal{C}$} is MDS if and only if any $k$ columns of $G$ are linearly independent.
Lemma 11.
( [51, Lemma 2.3]) Let $m$ be a fixed positive integer and let $I_{m}=\left\{0,\ldots,m-1\right\}=\left\{t_{1},\ldots,t_{s}\right\}\cup\left\{r_{1},\ldots,r_{s^{\prime}}\right\}$ be any partition of $I_{m}$ with $m=s+s^{\prime}$, $0=t_{1}<\ldots<t_{s}=m-1$ and $r_{1}<\ldots<r_{s^{\prime}}$. Then for any ${\mathcal{S}}=\left\{a_{1},\ldots,a_{s}\right\}\subseteq\mathbb{F}_{q}$, we have
$\det\left(\begin{array}[]{ccc}a_{1}^{t_{1}}&\ldots&a_{s}^{t_{1}}\\ a_{1}^{t_{2}}&\ldots&a_{s}^{t_{2}}\\ \vdots&\ddots&\vdots\\ a^{t_{s}}_{1}&\ldots&a^{t_{s}}_{s}\\ \end{array}\right)=\prod_{1\leq i<j\leq s}(a_{j}-a_{i})\cdot\det\left(\begin{array}[]{ccc}\sigma_{s-r_{1}}({\mathcal{S}})&\ldots&\sigma_{s-r_{s^{\prime}}}({\mathcal{S}})\\ \sigma_{s-r_{1}+1}({\mathcal{S}})&\ldots&\sigma_{s-r_{s^{\prime}}+1}({\mathcal{S}})\\ \vdots&\ddots&\vdots\\ \sigma_{s-r_{1}+s^{\prime}-1}({\mathcal{S}})&\ldots&\sigma_{s-r_{s^{\prime}}+s^{\prime}-1}({\mathcal{S}})\\ \end{array}\right).$
III ECPs of extended Han-Zhang codes
From now on, we explore the answers to Problem 1. In this section, we study the existence and specific forms of ECPs of extended Han-Zhang codes, and further provide a decoding algorithm for extended Han-Zhang codes based on their ECPs.
III-A The existence of $\ell$-ECPs of extended Han-Zhang codes
First of all, we note that the following nested relationships between extended Han-Zhang codes and (extended) GRS codes.
Lemma 12.
For any $3\leq k\leq k^{\prime}-1\leq n-2\leq q-3$ and $\gamma\in\mathbb{F}_{q}\setminus{\mathcal{S}}$, we have
${\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)\subseteq{\mathrm{GRS}}_{k+1}({\mathcal{S}},{{\mathbf{v}}},\infty)\subseteq{\mathrm{GRS}}_{k^{\prime}}({\mathcal{S}}^{\prime},{{\mathbf{v}}}^{\prime}),$
where ${\mathcal{S}}^{\prime}=\{(a_{1}-\gamma)^{-1},\ldots,(a_{n}-\gamma)^{-1},0\}\subseteq\mathbb{F}_{q}$ and ${{\mathbf{v}}}^{\prime}=(v_{1}(a_{1}-\gamma)^{k},\ldots,v_{n}(a_{n}-\gamma)^{k},1)\in(\mathbb{F}_{q}^{*})^{n+1}$.
Proof.
With (1) and (10), it is easily seen that
${\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)\subseteq{\mathrm{GRS}}_{k+1}({\mathcal{S}},{{\mathbf{v}}},\infty).$ (11)
Additionally, for any $\gamma\in\mathbb{F}_{q}\setminus{\mathcal{S}}$ (such $\gamma$ exists since $n\leq q-1$), it follows from [37] and [56, Lemma 4] that there exists a set ${\mathcal{S}}^{\prime}=\{a_{1}^{\prime},\ldots,a_{n+1}^{\prime}\}\subseteq\mathbb{F}_{q}$ and a vector ${{\mathbf{v}}}^{\prime}=(v_{1}^{\prime},\ldots,v_{n+1}^{\prime})\in(\mathbb{F}_{q}^{*})^{n+1}$ such that
${\mathrm{GRS}}_{k+1}({\mathcal{S}},{{\mathbf{v}}},\infty)={\mathrm{GRS}}_{k+1}({\mathcal{S}}^{\prime},{{\mathbf{v}}}^{\prime})\subseteq{\mathrm{GRS}}_{k^{\prime}}({\mathcal{S}}^{\prime},{{\mathbf{v}}}^{\prime})$ (12)
with
$a_{i}^{\prime}=\left\{\begin{array}[]{ll}(a_{i}-\gamma)^{-1},&{\rm if}~1\leq i\leq n,\\ 0,&{\rm if}~i=n+1,\end{array}\right.~{\rm and}~v_{i}^{\prime}=\left\{\begin{array}[]{ll}v_{i}(a_{i}-\gamma)^{k},&{\rm if}~1\leq i\leq n,\\ 1,&{\rm if}~i=n+1.\end{array}\right$.
Then the desired result immediately follows by combining (11) and (12). ∎
In the following, we investigate the existence of $\ell$-ECPs of extended Han-Zhang codes and give explicit forms of these $\ell$-ECPs if they exist.
Theorem 13.
Suppose that $\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor$. If $\mathcal{S}$} is $k$-zero-sum free, then the following statements hold.
  1. 1)
    If $2\nmid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ does not have an $\ell$-ECP. Moreover, ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $(\ell-1)$-ECP $(\mathcal{A},\mathcal{B})$ with
    $\mathcal{A}={\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf 1})~{\rm and}~\mathcal{B}={\mathrm{GRS}}_{\ell-1}({\mathcal{S}}^{\prime},{\bf v}^{\prime-1}\star{{\mathbf{u}}}^{\prime}),$
    where ${\mathcal{S}}^{\prime}=\{(a_{1}-\gamma)^{-1},\ldots,(a_{n}-\gamma)^{-1},0\}\subseteq\mathbb{F}_{q}$, ${{\mathbf{v}}}^{\prime}=(v_{1}(a_{1}-\gamma)^{k},\ldots,v_{n}(a_{n}-\gamma)^{k},1)\in(\mathbb{F}_{q}^{*})^{n+1}$ and ${{\mathbf{u}}}^{\prime}=(u_{1}^{\prime},\ldots,u_{n+1}^{\prime})\in(\mathbb{F}_{q}^{*})^{n+1}$ with
[…]
    for any $\gamma\in\mathbb{F}_{q}\setminus{\mathcal{S}}$.
  2. 2)
    If $2\mid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ with
    $\mathcal{A}={\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty)~{\rm and}~\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1).$
Proof.
1) Since $\mathcal{S}$} is $k$-zero-sum free, it follows from Lemma 4.1) that $d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+2$. From $2\nmid(n-k)$ and $3\leq k\leq n-2$, we get that $3\leq k\leq n-3$ and $2\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k+1}{2}\leq\frac{n-2}{2}<\frac{n+1}{2}-1$. Combining Lemmas 4.1) and 5, we immediately get that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)={\mathcal{C}}_{n-2\ell+1}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,n-2\ell+1,2\ell+1]_{q}$ non-GRS MDS code. By Lemma 7, if ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP, then it is a GRS code, a contradiction. Hence, ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ does not have an $\ell$-ECP.
Moreover, when ${\mathcal{S}}=\mathbb{F}_{q}$, it follows from [13, Corollary 1] that $\mathcal{S}$} is $k$-zero-sum free if and only if $q$ is even and k=q2k=q-2, which contradicts the fact $3\leq k\leq n-3$. Therefore, we conclude that $n\leq q-1$ and $\mathbb{F}_{q}\setminus{\mathcal{S}}\neq\emptyset$ under the given conditions. It then follows from Lemma 12 that
${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)={\mathcal{C}}_{n-2\ell+1}({\mathcal{S}},{\bf v},\infty)\subseteq{\mathrm{GRS}}_{n-2\ell+2}({\mathcal{S}},{\bf v},\infty)\subseteq{\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime}),$ (14)
where ${\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime})$ is an $n+1,n-2\ell+3,2\ell-1]_{q}$ GRS code with ${\mathcal{S}}^{\prime}=\{(a_{1}-\gamma)^{-1},\ldots,(a_{n}-\gamma)^{-1},0\}\subseteq\mathbb{F}_{q}$, ${{\mathbf{v}}}^{\prime}=(v_{1}(a_{1}-\gamma)^{k},\ldots,v_{n}(a_{n}-\gamma)^{k},1)\in(\mathbb{F}_{q}^{*})^{n+1}$, and any $\gamma\in\mathbb{F}_{q}\setminus{\mathcal{S}}$. Note that $2\ell-1=2(\ell-1)+1$ and $1\leq\ell-1<\frac{n+1}{2}-1$, then it follows from Lemmas 1, 7, and 8 that ${\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime})$ has an $(\ell-1)$-ECP $(\mathcal{A}^{\prime},\mathcal{B}^{\prime})$, where $\mathcal{A}^{\prime}={\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf r})$ is an $n+1,\ell,n-\ell+2]_{q}$ MDS code for some ${\bf r}\in(\mathbb{F}_{q}^{*})^{n+1}$ and
$\mathcal{B}^{\prime}=({\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf r})\star{\mathrm{GRS}}_{n-2\ell+3}({\mathcal{S}}^{\prime},{\bf v}^{\prime}))^{\perp}={\mathrm{GRS}}_{n-\ell+2}({\mathcal{S}}^{\prime},{\bf r}\star{\bf v}^{\prime})^{\perp}={\mathrm{GRS}}_{\ell-1}({\mathcal{S}}^{\prime},{\bf r}^{-1}\star{\bf v}^{-1}\star{{\mathbf{u}}}^{\prime})$
with ${{\mathbf{u}}}^{\prime}=(u_{1}^{\prime},\ldots,u_{n+1}^{\prime})$ and
$u_{i}^{\prime}=$ $\prod_{1\leq j\neq i\leq n+1}(a_{i}^{\prime}-a_{j}^{\prime})^{-1}$
$=$ $\left\{\begin{array}[]{ll}(a_{i}^{\prime}-a_{n+1}^{\prime})^{-1}\prod_{1\leq j\neq i\leq n}(a_{i}^{\prime}-a_{j}^{\prime})^{-1},&{\rm if}~1\leq i\leq n,\\ \prod_{1\leq j\leq n}(a_{n+1}^{\prime}-a_{j}^{\prime})^{-1},&{\rm if}~i=n+1,\end{array}\right$.
$=$ $\left\{\begin{array}[]{ll}(a_{i}-\gamma)^{n}\prod_{1\leq j\neq i\leq n}(a_{j}-\gamma)(a_{j}-a_{i})^{-1},&{\rm if}~1\leq i\leq n,\\ (-1)^{n}\prod_{1\leq j\leq n}(a_{j}-\gamma),&{\rm if}~i=n+1.\end{array}\right$.
Taking $\mathcal{A}={\bf r}^{-1}\star\mathcal{A}^{\prime}$ and $\mathcal{B}={\bf r}\star\mathcal{B}^{\prime}$, it then turns out from Lemma 6 and (14) that
$(\mathcal{A},\mathcal{B})=\left({\bf r}^{-1}\star\mathcal{A}^{\prime},{\bf r}\star\mathcal{B}^{\prime}\right)=\left({\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{\bf 1}),{\mathrm{GRS}}_{\ell-1}({\mathcal{S}}^{\prime},{\bf v}^{\prime-1}\star{{\mathbf{u}}}^{\prime})\right)$
is an $(\ell-1)$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This completes the proof of 1).
2) In this case, we have that $1\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k}{2}\leq\frac{n-3}{2}<\frac{n+1}{2}-1$ and ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has parameters $n+1,n-2\ell,2\ell+2]_{q}$. Since a parity-check matrix of a linear code is a generator matrix of its dual code, it then follows from Lemma 3 that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)^{\perp}$ has a generator matrix of the form
$\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{2\ell-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{2\ell-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{2\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{2\ell-1}&-1\\ \frac{u_{1}}{v_{1}}a_{1}^{2\ell}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{2\ell}&-\sum_{i=1}^{n}a_{i}\end{pmatrix}$. (15)
Take $\mathcal{A}={\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty)$ and $\mathcal{B}={\mathrm{GRS}}_{n+1-\ell}({\mathcal{S}},{\bf v},\infty)^{\perp}$. It follows from Lemma 1.2) and (1) that $\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)$, and $\mathcal{A}$ and $\mathcal{B}$ have generator matrices of the forms
$G_{\mathcal{A}}=\begin{pmatrix}1&\ldots&1&0\\ a_{1}&\ldots&a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ a^{\ell-1}_{1}&\ldots&a^{\ell-1}_{n}&0\\ a^{\ell}_{1}&\ldots&a^{\ell}_{n}&1\end{pmatrix}~{\rm and}~G_{\mathcal{B}}=\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-1}&-1\\ \end{pmatrix}$,
respectively.
It turns out that $\mathcal{A}\star\mathcal{B}$ is an $n+1,2\ell]_{q}$ linear code with generator matrix of the form
$G_{\mathcal{A}\star\mathcal{B}}=\left(\begin{array}[]{ccccc}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{2\ell-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{2\ell-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{2\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{2\ell-1}&-1\\ \end{array}\right),$
which, by combining with (15), implies that $\mathcal{A}\star\mathcal{B}\subseteq{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)^{\perp}$, $i.e.,$ the condition i) in Definition 5 holds. In addition, it can be checked that
$\begin{split}d(\mathcal{B}^{\perp})&=d({\mathrm{GRS}}_{n+1-\ell}({\mathcal{S}},{\bf v},\infty))=n+1-(n+1-\ell)+1=\ell+1>\ell,\\ \dim(\mathcal{A})&=\ell+1>\ell,~{\rm and}\\ d(\mathcal{A})+d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))&=n+1-(\ell+1)+1+2\ell+2=n+\ell+3>n+1,\end{split}$
which implies that the conditions ii), iii) and iv) in Definition 5 also hold. With Definition 5, we immediately conclude that
$(\mathcal{A},\mathcal{B})=\left({\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty),{\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)\right)$
is an $\ell$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This completes the proof of 2). ∎
Theorem 14.
Suppose that $\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor$. If $\mathcal{S}$} contains a $k$-zero-sum subset, then the following statements hold.
  1. 1)
    If $2\nmid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ with
    $\mathcal{A}=\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\ell+1}\}~{\rm and}~\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty).$
  2. 2)
    If $2\mid(n-k)$, then ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has an $\ell$-ECP $(\mathcal{A},\mathcal{B})$ with
    $\mathcal{A}={\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty)~{\rm and}~\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1).$
Proof.
1) Since $\mathcal{S}$} contains a $k$-zero-sum subset, it follows from Lemma 4.2) that $d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+1$. From $2\nmid(n-k)$ and $3\leq k\leq n-2$, we get that $3\leq k\leq n-3$ and $1\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k-1}{2}\leq\frac{n-4}{2}<\frac{n+1}{2}-1$ and ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has parameters $n+1,n-2\ell-1,2\ell+2]_{q}$. In this case, we derive from Lemma 3 that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)^{\perp}$ has a generator matrix of the form
$\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{2\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{2\ell-1}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{2\ell}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{2\ell}&-1\\ \frac{u_{1}}{v_{1}}a_{1}^{2\ell+1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{2\ell+1}&-\sum_{i=1}^{n}a_{i}\end{pmatrix}$. (16)
Take $\mathcal{A}=\{(c_{1},\ldots,c_{n},0):~(c_{1},\ldots,c_{n})\in{\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1})\}=\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\ell+1}\}$ and $\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty)$. According to (1), we deduce that $\mathcal{A}$ and $\mathcal{B}$ have generator matrices of the forms
$G_{\mathcal{A}}=\begin{pmatrix}1&\ldots&1&0\\ a_{1}&\ldots&a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ a^{\ell}_{1}&\ldots&a^{\ell}_{n}&0\end{pmatrix}~{\rm and}~G_{\mathcal{B}}=\begin{pmatrix}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-2}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-2}&0\\ \frac{u_{1}}{v_{1}}a_{1}^{\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{\ell-1}&1\\ \end{pmatrix}$,
respectively.
It is not difficult to show that $\mathcal{A}\star\mathcal{B}$ is an $n+1,2\ell]_{q}$ linear code with generator matrix of the form
$G_{\mathcal{A}\star\mathcal{B}}=\left(\begin{array}[]{cccc}\frac{u_{1}}{v_{1}}&\ldots&\frac{u_{n}}{v_{n}}&0\\ \frac{u_{1}}{v_{1}}a_{1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ \frac{u_{1}}{v_{1}}a_{1}^{2\ell-1}&\ldots&\frac{u_{n}}{v_{n}}a_{n}^{2\ell-1}&0\\ \end{array}\right).$
Combining with (16), we get that $\mathcal{A}\star\mathcal{B}\subseteq{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)^{\perp}$ and hence, the condition i) in Definition 5 holds. Moreover, since the dual code of an MDS code is still MDS, we can also verify that
$\begin{split}d(\mathcal{B}^{\perp})&=d({\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty)^{\perp})=\ell+1>\ell,\\ \dim(\mathcal{A})&=\ell+1>\ell,~{\rm and}\\ d(\mathcal{A})+d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))&=n+1-(\ell+1)+2\ell+2=n+\ell+2>n+1.\end{split}$
Therefore, the conditions ii), iii) and iv) in Definition 5 also hold. With Definition 5, we immediately conclude that
$(\mathcal{A},\mathcal{B})=\left(\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\ell+1}\},{\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty)\right)$
is an $\ell$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This completes the proof of 1).
2) In this case, we have that $1\leq\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor=\frac{n-k}{2}\leq\frac{n-3}{2}<\frac{n+1}{2}-1$ and ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ has parameters $n+1,n-2\ell,2\ell+1]_{q}$. Take $\mathcal{A}={\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty)$ and $\mathcal{B}={\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)$ as Theorem 13.2). By arguments similar to those in the proofs of Theorem 13.2) and 1) above, it suffices to show that
$d(\mathcal{A})+d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n+1-(\ell+1)+1+2\ell+1=n+\ell+2>n+1.$
As a result, we conclude that
$(\mathcal{A},\mathcal{B})=\left({\mathrm{GRS}}_{\ell+1}({\mathcal{S}},{\bf 1},\infty),{\mathrm{GRS}}_{\ell}({\mathcal{S}},{\bf v}^{-1}\star{{\mathbf{u}}},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)\right)$
is an $\ell$-ECP of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ by Definition 5. This completes the proof of 2). ∎
Remark 1.
Note that we suppose $\ell=\left\lfloor\frac{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))-1}{2}\right\rfloor$ for a more convenient representation in Theorems 13 and 14. Combining with Lemma 4, we can further deduce the explicit values of $\ell$ as listed in Table I, where the notation is the same with that used in Theorems 13 and 14.
TABLE I: $\ell$-ECPs $(\mathcal{A},\mathcal{B})$ of extended Han-Zhang codes ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$
${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ $2\mid(n-k)?$ $\ell$ ECP $(\mathcal{A},\mathcal{B})$ Result
$n+1,k,n-k+2]_{q}$ No $\frac{n-k-1}{2}$ $\left({\mathrm{GRS}}_{\frac{n-k+1}{2}}({\mathcal{S}}^{\prime},{\bf 1}),{\mathrm{GRS}}_{\frac{n-k-1}{2}}({\mathcal{S}}^{\prime},{\bf v}^{\prime-1}\star{\bf u}^{\prime})\right)$ Theorem 13.1)
$n+1,k,n-k+2]_{q}$ Yes $\frac{n-k}{2}$ $\left({\mathrm{GRS}}_{\frac{n-k+2}{2}}({\mathcal{S}},{\bf 1},\infty),{\mathrm{GRS}}_{\frac{n-k}{2}}({\mathcal{S}},{\bf v}^{-1}\star{\bf u},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)\right)$ Theorem 13.2)
$n+1,k,n-k+1]_{q}$ No $\frac{n-k-1}{2}$ $\left(\{(f(a_{1}),f(a_{2}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{\frac{n-k+1}{2}}\},{\mathrm{GRS}}_{\frac{n-k-1}{2}}({\mathcal{S}},{\bf v}^{-1}\star{\bf u},\infty)\right)$ Theorem 14.1)
$n+1,k,n-k+1]_{q}$ Yes $\frac{n-k}{2}$ $\left({\mathrm{GRS}}_{\frac{n-k+2}{2}}({\mathcal{S}},{\bf 1},\infty),{\mathrm{GRS}}_{\frac{n-k}{2}}({\mathcal{S}},{\bf v}^{-1}\star{\bf u},\infty){\mathrm{diag}}(\underbrace{1,\ldots,1}_{n},-1)\right)$ Theorem 14.2)
III-B An explicit decoding algorithm of extended Han-Zhang codes based on their $\ell$-ECPs
In this subsection, we provide an explicit decoding algorithm for extended Han-Zhang codes ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ based on their $\ell$-ECPs. First of all, we recall the error-correcting mechanism of $\ell$-ECPs for a given $n,k]_{q}$ linear code $\mathcal{C}$} (see also [16, 8]). Let $(\mathcal{A},\mathcal{B})$ be an $\ell$-ECP of $\mathcal{C}$} and let ${\bf y}={\bf c}+{\bf e}\in\mathbb{F}_{q}^{n}$ be a received vector with ${\bf c}\in{\mathcal{C}}$ and ${{\rm{wt}}}({\bf e})\leq\ell$. Note that based on conditions $\rm i)$$\rm iii)$} in Definition 5, Duursma and Kötter showed in [10, Theorem 1] that there exists a nonzero vector ${\bf a}\in\mathcal{A}$ such that
$\langle({\bf a}\star{\bf b}),{\bf y}\rangle=0,~\forall~{\bf b}\in\mathcal{B}$
or equivalently,
$G_{\mathcal{B}}\cdot{\mathrm{diag}}({\bf y})\cdot{\bf a}^{T}={\bf 0},$ (17)
where $G_{\mathcal{B}}$ denotes a generator matrix of $\mathcal{B}$ and ${\mathrm{diag}}({\bf y})$ represents the $n\times n$ diagonal matrix that has the elements of $\bf y$} on its main diagonal, and any solution ${\bf a}\in\mathcal{A}$ satisfies ${\bf e}\star{\bf a}={\bf 0}$. Then it is easily seen that ${\mathrm{supp}}({\bf e})\subseteq z({\bf a})$, and hence, any nonzero solution $\bf a$} can efficiently locate the possible error positions of $\bf y$}.
In addition, Pellikaan pointed out in [36] that the condition $\rm iv)$} $d(\mathcal{A})+d({\mathcal{C}})>n$” in Definition 5 can guarantee that
$H_{\mathcal{C}}{\bf x}^{T}=H_{\mathcal{C}}{\bf y}^{T}=H_{\mathcal{C}}{\bf e}^{T}$ (18)
has a unique solution, where $H_{{\mathcal{C}}}$ is a parity-check matrix of $\mathcal{C}$}. It then follows that ${\bf x}={\bf e}$ and $x_{i}=0$ for all $i\notin z({\bf a})$ if we write ${\bf x}=(x_{1},\ldots,x_{n})$. As a result, the $\ell$-ECP $(\mathcal{A},\mathcal{B})$ has decoded the received vector $\bf y$} to the codeword ${\bf c}={\bf y}-{\bf x}$.
Based on the above discussion, we conclude that to decode a linear code $\mathcal{C}$} using its $\ell$-ECP $(\mathcal{A},\mathcal{B})$, one should first determine explicit representations of the code pair $\mathcal{A}$ and $\mathcal{B}$, and then solve the system of equations (17) and (18). By combining Theorems 13 and 14, we can now present a decoding algorithm for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ based on its $\ell$-ECP as follows.
Theorem 15.
If $\mathcal{S}$} is $k$-zero-sum free and $2\nmid(n-k)$, then an extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ can be decoded by Algorithm 1 in $O(n^{3})$ time. Moreover, the decoding algorithms with the same time complexity for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ with respect to the other three cases discussed in Theorems 13 and 14 can be similarly derived by modifying Lines 8-14 of Algorithm 1 based on Table I accordingly.
Proof.
The decoding correctness is straightforward by combining Theorems 13 and 14 as well as the discussions above. By [36, Theorem 2.14], the time complexity is $O(n^{3})$. More details can be found in Algorithm 1. ∎
Input: A received vector ${\bf y}=(y_{1},\ldots,y_{n+1})\in\mathbb{F}_{q}^{n+1}$
Output: A decoded codeword ${\bf c}\in{\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)$ or a failure message
1 begin
2 $H_{{\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)}\leftarrow$ the parity-check matrix of ${\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)$ given by (10)
3 if $H_{{\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)}\cdot{\bf y}^{T}={\bf 0}$ then
4    ${\bf c}\leftarrow{\bf y}$
5    return $\bf c$} is the decoded codeword
6 else
7    ${\mathrm{diag}}({\bf y})\leftarrow{\mathrm{diag}}(y_{1},\ldots,y_{n+1})$
8    $\ell\leftarrow\frac{n-k-1}{2}$
9    $\gamma\leftarrow$ arbitrary nonzero element of $\mathbb{F}_{q}\setminus{\mathcal{S}}$
10    ${\mathcal{S}}^{\prime}\leftarrow\{(a_{1}-\gamma)^{-1},\ldots,(a_{n}-\gamma)^{-1},0\}$
11    ${{\mathbf{v}}}^{\prime}\leftarrow(v_{1}(a_{1}-\gamma)^{k-1},\ldots,v_{n}(a_{n}-\gamma)^{k-1},1)$
12    ${{\mathbf{u}}}^{\prime}\leftarrow(u_{1}^{\prime},\ldots,u_{n+1}^{\prime})$ with $u_{i}^{\prime}$ given in (13)
13    $G_{\mathcal{A}}\leftarrow$ the generator matrix of ${\mathrm{GRS}}_{\ell+1}({\mathcal{S}}^{\prime},{\bf 1})$ given by (1)
14    $G_{\mathcal{B}}\leftarrow$ the generator matrix of ${\mathrm{GRS}}_{\ell}({\mathcal{S}}^{\prime},{{\mathbf{v}}}^{\prime-1}\star{{\mathbf{u}}}^{\prime})$ given by (1)
15    ${\bf s}({\bf y})\leftarrow\{{\bf s}=(s_{1},\ldots,s_{\dim(\mathcal{A})}):~G_{\mathcal{B}}\cdot{\mathrm{diag}}({\bf y})\cdot G_{\mathcal{A}}^{T}\cdot{\bf s}^{T}={\bf 0}\}$
16
17    if ${\bf s}({\bf y})=\{{\bf 0}\}$ then
18       return $\bf y$} has more than $\ell$ errors
19    else
20       ${\bf s}_{0}\leftarrow$ any nonzero vector in ${\bf s}({\bf y})$
21       ${\bf a}\leftarrow{\bf s}_{0}G_{\mathcal{A}}$
22       $Z\leftarrow z({\bf a})$
23       ${\bf x}({\bf y})\leftarrow\{{\bf x}:~H_{{\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)}\cdot{\bf x}^{T}=H_{{\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)}\cdot{\bf y}^{T}~{\rm with}~x_{i}=0,~\forall~i\notin Z\}$
24
25       if ${\bf x}({\bf y})=\{{\bf x}_{0}\}$ and ${{\rm{wt}}}({\bf x}_{0})\leq\ell$ then
26          ${\bf c}\leftarrow{\bf y}-{\bf x}_{0}$
27          return $\bf c$} is the decoded codeword
28       else
29          return $\bf y$} has more than $\ell$ errors
30         end if
31
32      end if
33
34   end if
35
36 end
37
Algorithm 1 An algorithm for decoding ${\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)$ when $\mathcal{S}$} is $k$-zero-sum free and $2\nmid(n-k)$
Finally, we give some specific examples to illustrate the decoding procedure given in Algorithm 1.
Example 1.
Let $q=17$, ${\mathcal{S}}=\{1,3,5,7,10,12,14,16\}\subseteq\mathbb{F}_{17}$, and ${{\mathbf{v}}}=(1,1,1,1,1,1,1,1)\in(\mathbb{F}_{17}^{*})^{8}$. Consider the extended Han-Zhang code ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ with generator matrix
$G_{3}=\begin{pmatrix}1&1&1&1&1&1&1&1&0\\ 1&3&5&7&10&12&14&16&0\\ 1&10&6&3&14&11&7&16&1\end{pmatrix}$.
Suppose that the received vector is ${\bf y}=(4,6,1,14,5,7,12,15,2)\in\mathbb{F}_{17}^{9}$ and we have
${\mathrm{diag}}({\bf y})={\mathrm{diag}}(4,6,1,14,5,7,12,15,2).$
We have the following results.
  • Note that $\mathcal{S}$} is $3$-zero-sum free. It then follows from Lemmas 3, 4.1), and 5 that the extended Han-Zhang code ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ is a $9,3,7]_{17}$ non-GRS MDS code with parity-check matrix
    $H_{{\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)}=\begin{pmatrix}4&1&11&13&4&6&16&13&0\\ 4&3&4&6&6&4&3&4&0\\ 4&9&3&8&9&14&8&13&0\\ 4&10&15&5&5&15&10&4&0\\ 4&13&7&1&16&10&4&13&16\\ 4&5&1&7&7&1&5&4&0\end{pmatrix}$.
    Note also that
    $H_{{\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)}\cdot{\bf y}^{T}=(1,4,7,12,13,1)^{T}\neq{\bf 0},$
    so $\bf y$} is not a codeword of ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$.
  • From Lines 8 to 14 of Algorithm 1, we have $\ell=2$ and we can take $\gamma=2\in\mathbb{F}_{17}\setminus{\mathcal{S}}$, ${\mathcal{S}}^{\prime}=\{16,1,6,7,15,12,10,11,0\}$, ${{\mathbf{v}}}^{\prime}=(1,1,9,8,13,15,8,9,1)$, and ${{\mathbf{u}}}^{\prime}=(12,14,11,1,7,12,4,4,3)$. Then $(\mathcal{A},\mathcal{B})$ is a $2$-ECP of ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$, where
    $\mathcal{A}={\mathrm{GRS}}_{3}({\mathcal{S}}^{\prime},{\bf 1})~\mbox{and}~\mathcal{B}={\mathrm{GRS}}_{2}({\mathcal{S}}^{\prime},{\bf v}^{\prime-1}\star{\bf u}^{\prime}).$
    Moreover, we deduce that
    $G_{\mathcal{A}}=\begin{pmatrix}1&1&1&1&1&1&1&1&1\\ 16&1&6&7&15&12&10&11&0\\ 1&1&2&15&4&8&15&2&0\\ \end{pmatrix}~\mbox{and}~G_{\mathcal{B}}=\begin{pmatrix}5&14&13&3&12&13&5&3&3\\ 12&14&10&4&10&3&16&16&0\end{pmatrix}$.
  • With Line 15 of Algorithm 1, we solve the system of linear equations
    $G_{\mathcal{B}}\cdot{\mathrm{diag}}({\bf y})\cdot G_{\mathcal{A}}^{T}\cdot{\bf s}^{T}={\bf 0}$
    and get ${\bf s}({\bf y})=\{\mu(1,6,10):~\mu\in\mathbb{F}_{17}\}$. Take ${\bf s}_{0}=(1,6,10)$, then we have
    ${\bf a}={\bf s}_{0}G_{\mathcal{A}}=(5,0,6,6,12,0,7,2,1)~{\rm and}~Z=z({\bf a})=\{2,6\}.$
  • Based on Line 22 of Algorithm 1, we further set ${\bf x}=(0,x_{2},0,0,0,x_{6},0,0,0)$ and solve the system of linear equations
    $H_{{\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)}{\bf x}^{T}=H_{{\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)}{\bf y}^{T}.$
    As a result, we get the solution ${\bf x}_{0}=\{0,16,0,0,0,6,0,0,0\}$ and ${{\rm{wt}}}({\bf x}_{0})=2$.
  • Finally, according to Lines 24 and 25 of Algorithm 1, the received vector $\bf y$} is then decoded to the codeword
    ${\bf c}={\bf y}-{\bf x}_{0}=(4,7,1,14,5,1,12,15,2).$
    Note also that $H_{{\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)}\cdot{\bf c}^{T}={\bf 0}$. Hence, the decoding process is successful.
Example 2.
Let $q=16$ and let $\omega$} be a primitive element of $\mathbb{F}_{16}$ such that ${\omega}^{4}+{\omega}+1=0$, ${\mathcal{S}}=\{1,{\omega},{\omega}^{2},{\omega}^{3},{\omega}^{4},{\omega}^{6},{\omega}^{7},{\omega}^{8},{\omega}^{9},{\omega}^{11},\\ {\omega}^{12},{\omega}^{13},{\omega}^{14}\}\subseteq\mathbb{F}_{16}$, and ${{\mathbf{v}}}=(1,1,1,1,1,1,1,1,1,1,1,1,1)\in(\mathbb{F}_{16}^{*})^{13}$. Consider the extended Han-Zhang code ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$ with generator matrix
$G_{7}=\left(\begin{array}[]{cccccccccccccc}1&1&1&1&1&1&1&1&1&1&1&1&1&0\\ 1&{\omega}&{\omega}^{2}&{\omega}^{3}&{\omega}^{4}&{\omega}^{6}&{\omega}^{7}&{\omega}^{8}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{13}&{\omega}^{14}&0\\ 1&{\omega}^{2}&{\omega}^{4}&{\omega}^{6}&{\omega}^{8}&{\omega}^{12}&{\omega}^{14}&{\omega}&{\omega}^{3}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{13}&0\\ 1&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&0\\ 1&{\omega}^{4}&{\omega}^{8}&{\omega}^{12}&{\omega}&{\omega}^{9}&{\omega}^{13}&{\omega}^{2}&{\omega}^{6}&{\omega}^{14}&{\omega}^{3}&{\omega}^{7}&{\omega}^{11}&0\\ 1&{\omega}^{5}&{\omega}^{10}&1&{\omega}^{5}&1&{\omega}^{5}&{\omega}^{10}&1&{\omega}^{10}&1&{\omega}^{5}&{\omega}^{10}&0\\ 1&{\omega}^{7}&{\omega}^{14}&{\omega}^{6}&{\omega}^{13}&{\omega}^{12}&{\omega}^{4}&{\omega}^{11}&{\omega}^{3}&{\omega}^{2}&{\omega}^{9}&{\omega}&{\omega}^{8}&1\\ \end{array}\right).$
Suppose that the received vector is ${\bf y}=({\omega}^{12},{\omega},{\omega}^{13},{\omega}^{7},1,{\omega}^{3},{\omega}^{9},{\omega},{\omega}^{2},{\omega},{\omega}^{7},{\omega}^{6},{\omega}^{12},{\omega}^{11})\in\mathbb{F}_{16}^{14}$ and we have
${\mathrm{diag}}({\bf y})={\mathrm{diag}}({\omega}^{12},{\omega},{\omega}^{13},{\omega}^{7},1,{\omega}^{3},{\omega}^{9},{\omega},{\omega}^{2},{\omega},{\omega}^{7},{\omega}^{6},{\omega}^{12},{\omega}^{11}).$
We have the following results.
  • Note that $\mathcal{S}$} contains a $7$-zero-sum subset, such as the $7$-subset $\{1,{\omega}^{12},{\omega}^{2},{\omega}^{13},{\omega}^{3},{\omega}^{8},{\omega}^{9}\}$. It then follows from Lemmas 3 and 4.2) that the extended Han-Zhang code ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$ is a $14,7,7]_{16}$ NMDS code with parity-check matrix
    $H_{{\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)}=\left(\begin{array}[]{cccccccccccccc}1&{\omega}^{11}&{\omega}^{7}&{\omega}^{11}&{\omega}^{14}&{\omega}^{7}&{\omega}^{11}&{\omega}^{13}&{\omega}^{13}&{\omega}^{13}&{\omega}^{14}&{\omega}^{14}&{\omega}^{7}&0\\ 1&{\omega}^{12}&{\omega}^{9}&{\omega}^{14}&{\omega}^{3}&{\omega}^{13}&{\omega}^{3}&{\omega}^{6}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{6}&0\\ 1&{\omega}^{13}&{\omega}^{11}&{\omega}^{2}&{\omega}^{7}&{\omega}^{4}&{\omega}^{10}&{\omega}^{14}&{\omega}&{\omega}^{5}&{\omega}^{8}&{\omega}^{10}&{\omega}^{5}&0\\ 1&{\omega}^{14}&{\omega}^{13}&{\omega}^{5}&{\omega}^{11}&{\omega}^{10}&{\omega}^{2}&{\omega}^{7}&{\omega}^{10}&{\omega}&{\omega}^{5}&{\omega}^{8}&{\omega}^{4}&0\\ 1&1&1&{\omega}^{8}&1&{\omega}&{\omega}^{9}&1&{\omega}^{4}&{\omega}^{12}&{\omega}^{2}&{\omega}^{6}&{\omega}^{3}&0\\ 1&{\omega}&{\omega}^{2}&{\omega}^{11}&{\omega}^{4}&{\omega}^{7}&{\omega}&{\omega}^{8}&{\omega}^{13}&{\omega}^{8}&{\omega}^{14}&{\omega}^{4}&{\omega}^{2}&1\\ 1&{\omega}^{2}&{\omega}^{4}&{\omega}^{14}&{\omega}^{8}&{\omega}^{13}&{\omega}^{8}&{\omega}&{\omega}^{7}&{\omega}^{4}&{\omega}^{11}&{\omega}^{2}&{\omega}&1\end{array}\right).$
    Note also that
    $H_{{\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)}\cdot{\bf y}^{T}=({\omega}^{11},{\omega}^{11},{\omega}^{10},{\omega}^{3},{\omega}^{6},{\omega}^{6},{\omega}^{9})^{T}\neq{\bf 0},$
    so $\bf y$} is not a codeword of ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$.
  • Modifying Lines 8 to 14 of Algorithm 1 based on Table I, we have $\ell=3$ and ${{\mathbf{u}}}=(1,{\omega}^{11},{\omega}^{7},{\omega}^{11},{\omega}^{14},{\omega}^{7},{\omega}^{11},{\omega}^{13},{\omega}^{13},{\omega}^{13},\\ {\omega}^{14},{\omega}^{14},{\omega}^{7})$. Then $(\mathcal{A},\mathcal{B})$ is a $3$-ECP of ${\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)$, where
    $\mathcal{A}=\{(f(a_{1}),\ldots,f(a_{n}),0):~f(x)\in\mathbb{F}_{q}[x]_{4}\}~\mbox{and}~\mathcal{B}={\mathrm{GRS}}_{3}({\mathcal{S}},{\bf v}^{-1}\star{\bf u},\infty).$
    Moreover, we deduce that
    $G_{\mathcal{A}}=\left(\begin{array}[]{cccccccccccccc}1&1&1&1&1&1&1&1&1&1&1&1&1&0\\ 1&{\omega}&{\omega}^{2}&{\omega}^{3}&{\omega}^{4}&{\omega}^{6}&{\omega}^{7}&{\omega}^{8}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{13}&{\omega}^{14}&0\\ 1&{\omega}^{2}&{\omega}^{4}&{\omega}^{6}&{\omega}^{8}&{\omega}^{12}&{\omega}^{14}&{\omega}&{\omega}^{3}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{13}&0\\ 1&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&{\omega}^{3}&{\omega}^{6}&{\omega}^{9}&{\omega}^{12}&0\\ \end{array}\right)~\mbox{and}$~
    $G_{\mathcal{B}}=\left(\begin{array}[]{cccccccccccccc}1&1&1&1&1&1&1&1&1&1&1&1&1&0\\ 1&{\omega}^{12}&{\omega}^{9}&{\omega}^{14}&{\omega}^{3}&{\omega}^{13}&{\omega}^{3}&{\omega}^{6}&{\omega}^{7}&{\omega}^{9}&{\omega}^{11}&{\omega}^{12}&{\omega}^{6}&0\\ 1&{\omega}^{13}&{\omega}^{11}&{\omega}^{2}&{\omega}^{7}&{\omega}^{4}&{\omega}^{10}&{\omega}^{14}&{\omega}&{\omega}^{5}&{\omega}^{8}&{\omega}^{10}&{\omega}^{5}&1\\ \end{array}\right).$
  • With Line 15 of Algorithm 1, we solve the system of linear equations
    $G_{\mathcal{B}}\cdot{\mathrm{diag}}({\bf y})\cdot G_{\mathcal{A}}^{T}\cdot{\bf s}^{T}={\bf 0}$
    and get ${\bf s}({\bf y})=\{\mu(1,1,{\omega}^{8},{\omega}^{4}):~\mu\in\mathbb{F}_{16}\}$. Take ${\bf s}_{0}=(1,1,{\omega}^{8},{\omega}^{4})$, then we have
    ${\bf a}={\bf s}_{0}G_{\mathcal{A}}=({\omega}^{5},{\omega}^{12},{\omega}^{13},{\omega}^{13},{\omega},0,{\omega}^{5},{\omega}^{4},{\omega}^{10},{\omega}^{8},{\omega}^{13},{\omega},{\omega}^{5},0)~{\rm and}~Z=z({\bf a})=\{6,14\}.$
  • Based on Line 22 of Algorithm 1, we further set ${\bf x}=(0,0,0,0,0,x_{6},0,0,0,0,0,0,0,x_{14})$ and find that the system of linear equations
    $H_{{\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)}{\bf x}^{T}=H_{{\mathcal{C}}_{7}({\mathcal{S}},{{\mathbf{v}}},\infty)}{\bf y}^{T}$
    has no solution. According to Line 27 of Algorithm 1, we immediately conclude that the received vector $\bf y$} incurs more than 33 errors.
IV Deep holes of extended Han-Zhang codes
In addition to the existence of ECPs, deep holes of linear codes are also closely related to their maximum-likelihood decoding problems. In this section, we discuss the deep holes of the extended Han-Zhang codes and apply the deep holes to obtain more non-GRS MDS codes with larger lengths and dimensions. Moreover, we discuss their connections to the well-known Roth-Lempel codes. According to Definition 6, we first need to determine the covering radius of extended Han-Zhang codes.
Theorem 16.
For any extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$, we have
$\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+1.$
Proof.
With Definition 4, it follows from the redundancy bound ( [21, Corollary 11.1.3]) that
$\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\leq n-k+1.$ (19)
On the other hand, by Lemma 12, we have that ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)\subseteq{\mathrm{GRS}}_{k+1}({\mathcal{S}},{\bf v},\infty)$, where ${\mathrm{GRS}}_{k+1}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,k+1,n-k+1]_{q}$ MDS code. Combining with the supercode lemma ( [21, Corollary 11.1.5]), it then implies that
$\begin{split}\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\geq\min\{{{\rm{wt}}}(\mathbf{c}):\mathbf{c}\in{\mathrm{GRS}}_{k+1}({\mathcal{S}},{\bf v},\infty)\setminus{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)\}\geq n-k+1.\end{split}$ (20)
We complete the proof by considering (19) and (20) together. ∎
IV-A Determining deep holes of extended Han-Zhang codes
Zhuang etal.et~al. provided in [58, Proposition 2] an equivalent condition under which a vector is a deep hole of a Reed-Solomon code, utilizing the covering radius and a generator matrix of this code. More recently, Wu etal.et~al. extended this result to general MDS codes in [48, Lemma 5]. However, as shown in Lemma 4, the code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is not always an MDS code. Consequently, the results from [48, Lemma 5] and [58, Proposition 2] cannot be directly applied to determine the deep holes of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This limitation motivates the following theorem.
Theorem 17.
Suppose that $\mathbf{u}\in\mathbb{F}_{q}^{n+1}$ and $\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ \mathbf{u}\end{array}\right)$ generates a linear code $\mathcal{C}$$^{\prime}$. Then $\mathbf{u}$ is a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if $\mathcal{C}$$^{\prime}$ is an $n+1,k+1,n-k+1]_{q}$ MDS code.
Proof.
Note that if $\mathbf{u}$ is a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$, then $\mathbf{u}\notin{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$, which implies that $\mathcal{C}$$^{\prime}$ is an $n+1,k+1]_{q}$ linear code. Note that
$d({\mathcal{C}}^{\prime})=\min\{d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)),d(\mathbf{u},{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\}$
and it follows from Lemmas 4 and 16 as well as Definition 6 that
$d({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\geq n-k+1=\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))\geq d(\mathbf{u},{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)).$
Hence, $d({\mathcal{C}}^{\prime})=d(\mathbf{u},{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))$. It further implies that $\mathbf{u}$ is a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if $d({\mathcal{C}}^{\prime})=d(\mathbf{u},{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+1$, if and only if $\mathcal{C}$$^{\prime}$ is an $n+1,k+1,n-k+1]_{q}$ MDS code. This completes the proof. ∎
Lemma 18.
Suppose that $\mathbf{g}=(g(a_{1}),\ldots,g(a_{n}),u_{n+1})\in\mathbb{F}_{q}^{n+1}$, where $g(x)=g^{\prime}(x)+f(x)\in\mathbb{F}_{q}[x]$ with $f(x)\in\mathcal{V}_{k}$. If $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ for $v_{n+1}\in\mathbb{F}_{q}$, then $g^{\prime}(x)\in\mathbb{F}_{q}[x]\setminus\mathcal{V}_{k}$.
Proof.
Assume that $g^{\prime}(x)\in\mathcal{V}_{k}$. Then $g(x)=g^{\prime}(x)+f(x)\in\mathcal{V}_{k}$ and
$(\mathbf{v},v_{n+1})\star\mathbf{g}=(v_{1}(g^{\prime}(a_{1})+f(a_{1})),\ldots,v_{n}(g^{\prime}(a_{n})+f(a_{n})),u_{n+1}v_{n+1}).$
It can be checked that
$\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right)=\left(\begin{array}[]{ccccc}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1\\ v_{1}(g^{\prime}(a_{1})+f(a_{1}))&\ldots&v_{n}(g^{\prime}(a_{n})+f(a_{n}))&u_{n+1}v_{n+1}\end{array}\right)$
and it can be transformed to the matrix
$\left(\begin{array}[]{ccccc}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1\\ 0&\ldots&0&u_{n+1}v_{n+1}-f_{k}-g_{k}\end{array}\right)$ (27)
by a series of elementary row operations, where $f_{k}$ and $g_{k}$ are respective coefficients of $x^{k}$ in $f(x)$ and $g^{\prime}(x)$. From Definition 4, we have $k+1\leq n$. It then follows that the determinants of all the $(k+1)\times(k+1)$ matrices consisting of any $k+1$ columns of the first $n$ columns of the matrix given in (27) equal zero. With Lemma 10, we immediately conclude that the linear code generated by $\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right)$ is not MDS. According to Theorem 17, we get that $g^{\prime}(x)\in\mathbb{F}_{q}[x]\setminus\mathcal{V}_{k}$ if $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. This completes the proof. ∎
With Lemma 18, we can always assume that no term in $g^{\prime}(x)$ belongs to $\mathcal{V}_{k}$. In the following, we consider two cases of $g^{\prime}(x)$ in terms of their lowest degrees. In other words, we consider $g^{\prime}(x)=g_{k-1}x^{k-1}$ with $g_{k-1}\in\mathbb{F}_{q}^{*}$ and $g^{\prime}(x)=g_{k+1}x^{k+1}+g_{k-1}x^{k-1}$ with $g_{k+1}\in\mathbb{F}_{q}^{*}$ and $g_{k-1}\in\mathbb{F}_{q}$ in the following.
Theorem 19.
Suppose that $\mathbf{g}=(g(a_{1}),\ldots,g(a_{n}),u_{n+1})\in\mathbb{F}_{q}^{n+1}$, where $g(x)=g_{k-1}x^{k-1}+f(x)\in\mathbb{F}_{q}[x]$ with $g_{k-1}\in\mathbb{F}_{q}^{*}$ and $f(x)\in\mathcal{V}_{k}$. Then for any $v_{n+1}\in\mathbb{F}_{q}$, $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if one of the following conditions holds:
  1. 1)
    $u_{n+1}v_{n+1}=f_{k}$;
  2. 2)
    $u_{n+1}v_{n+1}\neq f_{k}$ and $\mathcal{S}$} is an $(n,k,g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1})$-set.
Proof.
Since $(\mathbf{v},v_{n+1})\star\mathbf{g}=(v_{1}g(a_{1}),\ldots,v_{n}g(a_{n}),u_{n+1}v_{n+1})$, we have that
$\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right)\thicksim\left(\begin{array}[]{ccccc}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0\\ v_{1}a_{1}^{k-1}&\ldots&v_{n}a_{n}^{k-1}&g_{k-1}^{-1}\left(u_{n+1}v_{n+1}-f_{k}\right)\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1\\ \end{array}\right),$ (36)
[…]
Case 1: $u_{n+1}v_{n+1}=f_{k}$. Then the linear code generated by the matrix given in (36) is an $n+1,k+1,n-k+1]_{q}$ MDS code, actually an EGRS code. According to Theorem 17, $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$.
Case 2: $u_{n+1}v_{n+1}\neq f_{k}$. We note that the linear code generated by the matrix given in (36) is equivalent to the linear code $\mathcal{C}$$^{\prime}$ generated by the matrix
$\left(\begin{array}[]{ccccc}1&\ldots&1&0\\ a_{1}&\ldots&a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ a_{1}^{k-2}&\ldots&a_{n}^{k-2}&0\\ a_{1}^{k-1}&\ldots&a_{n}^{k-1}&1\\ a_{1}^{k}&\ldots&a_{n}^{k}&g_{k-1}\left(u_{n+1}v_{n+1}-f_{k}\right)^{-1}\\ \end{array}\right).$
Note also that $\mathcal{C}$$^{\prime}$ coincides with a punctured code of the $n+2,k+1]_{q}$ Roth-Lempel code $\mathrm{RL}_{k+1,\delta}({\mathcal{S}})$ with $\delta=g_{k-1}\left(u_{n+1}v_{n+1}-f_{k}\right)^{-1}$, defined as in Definition 2, by deleting its penultimate coordinate. It then follows from Lemma 2 and the fact that a punctured code of an MDS code is still MDS that $\mathcal{C}$$^{\prime}$ is an $n+1,k+1,n-k+1]_{q}$ MDS code if $\mathcal{S}$} is an $\left(n,k,g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1}\right)$-set.
Conversely, if $\mathcal{C}$$^{\prime}$ is an $n+1,k+1,n-k+1]_{q}$ MDS code with generator matrix $G^{\prime}$, then the linear code generated by the matrix
$\overline{G^{\prime}}=\left(G^{\prime},~~(0,\ldots,0,1)^{T}\right),$
that is, the matrix obtained by appending a column vector $(0,\ldots,0,1)^{T}$ of length $k+1$ to $G^{\prime}$, is necessarily MDS and this linear code is just the Roth-Lempel code ${\mathrm{RL}}_{k+1,\delta}({\mathcal{S}})$, where $\delta=g_{k-1}\left(u_{n+1}v_{n+1}-f_{k}\right)^{-1}$. Again, from Lemma 2, we conclude that $\mathcal{S}$} is an $\left(n,k,g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1}\right)$-set. Hence, $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if $\mathcal{S}$} is an $(n,k,g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1})$-set.
Combining Cases 1 and 2, we complete the proof. ∎
Let $\sigma_{i}(x_{1},\ldots,x_{n})$ be the $i$th elementary symmetric function in $n$ variables $x_{1},\ldots,x_{n}$ such that
$\sigma_{i}(x_{1},\ldots,x_{n})=\left\{\begin{array}[]{ll}1,&{\rm if}~i=0,\\ \sum_{I\subseteq\{1,\ldots,n\},~|I|=i}\prod_{j\in I}x_{j},&{\rm if}~1\leq i\leq n,\\ 0,&{\rm if}~i<0~{\rm or}~i>n.\end{array}\right$.
Furthermore, for any set $\mathcal{T}=\{t_{1},\ldots,t_{n}\}\subseteq\mathbb{F}_{q}$, we let
$\sigma_{i}(\mathcal{T})=\sigma_{i}(t_{1},\ldots,t_{n}).$
With this notation, we have the following theorem.
Theorem 20.
Suppose that $\mathbf{g}=(g(a_{1}),\ldots,g(a_{n}),u_{n+1})\in\mathbb{F}_{q}^{n+1}$, where $g(x)=g_{k+1}x^{k+1}+g_{k-1}x^{k-1}+f(x)\in\mathbb{F}_{q}[x]$ with $g_{k+1}\in\mathbb{F}_{q}^{*}$, $g_{k-1}\in\mathbb{F}_{q}$, and $f(x)\in\mathcal{V}_{k}$. Then for any $v_{n+1}\in\mathbb{F}_{q}$, $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if
$g_{k-1}\notin\left\{g_{k+1}\left(\sigma_{2}({\mathcal{S}}_{k})-\sigma_{1}({\mathcal{S}}_{k})^{2}\right)+(u_{n+1}v_{n+1}-f_{k})\sigma_{1}({\mathcal{S}}_{k}):~{\mathcal{S}}_{k}\subseteq{\mathcal{S}}\right\}\cup\left\{g_{k+1}\sigma_{2}({\mathcal{S}}_{k+1}):~{\mathcal{S}}_{k+1}\subseteq{\mathcal{S}}\right\}.$
Proof.
Under the given condition, we know that $\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right)$ is row equivalent to the following matrix
$\left(\begin{array}[]{ccccc}v_{1}&\ldots&v_{n}&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1\\ v_{1}(g_{k+1}a_{1}^{k+1}+g_{k-1}a_{1}^{k-1})&\ldots&v_{n}(g_{k+1}a_{n}^{k+1}+g_{k-1}a_{n}^{k-1})&u_{n+1}v_{n+1}-f_{k}\\ \end{array}\right)$ (43)
Let $M_{k+1}$ be any $(k+1)\times(k+1)$ submatrix of the matrix given in (43). Since $k+1\leq n$, to determine whether the linear code generated by $\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right)$ is MDS, we need to consider the following two cases.
Case 1: $M_{k+1}$ contains the last column of the matrix given in (43). In this case, $M_{k+1}$ has the following form
$\left(\begin{array}[]{cccc}v_{i_{1}}&\ldots&v_{i_{k}}&0\\ v_{i_{1}}a_{i_{1}}&\ldots&v_{i_{k}}a_{i_{k}}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{i_{1}}a^{k-2}_{i_{1}}&\ldots&v_{i_{k}}a^{k-2}_{i_{k}}&0\\ v_{i_{1}}a^{k}_{i_{1}}&\ldots&v_{i_{k}}a^{k}_{i_{k}}&1\\ v_{i_{1}}(g_{k+1}a^{k+1}_{i_{1}}+g_{k-1}a^{k-1}_{i_{1}})&\ldots&v_{i_{k}}(g_{k+1}a^{k+1}_{i_{k}}+g_{k-1}a^{k-1}_{i_{k}})&u_{n+1}v_{n+1}-f_{k}\\ \end{array}\right),$
where ${\mathcal{S}}_{k}=\{a_{i_{1}},\ldots,a_{i_{k}}\}\subseteq{\mathcal{S}}$. It then follows from Lemma 11 that the determinant of $M_{k+1}$ is
$-g_{k-1}\det\left(\begin{array}[]{cccc}v_{i_{1}}&\ldots&v_{i_{k}}&0\\ v_{i_{1}}a_{i_{1}}&\ldots&v_{i_{k}}a_{i_{k}}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{i_{1}}a^{k-1}_{i_{1}}&\ldots&v_{i_{k}}a^{k-1}_{i_{k}}&0\\ v_{i_{1}}a^{k}_{i_{1}}&\ldots&v_{i_{k}}a^{k}_{i_{k}}&1\\ \end{array}\right)+\det\left(\begin{array}[]{cccc}v_{i_{1}}&\ldots&v_{i_{k}}&0\\ v_{i_{1}}a_{i_{1}}&\ldots&v_{i_{k}}a_{i_{k}}&0\\ \vdots&\ddots&\vdots&\vdots\\ v_{i_{1}}a^{k-2}_{i_{1}}&\ldots&v_{i_{k}}a^{k-2}_{i_{k}}&0\\ v_{i_{1}}a^{k}_{i_{1}}&\ldots&v_{i_{k}}a^{k}_{i_{k}}&1\\ v_{i_{1}}g_{k+1}a^{k+1}_{i_{1}}&\ldots&v_{i_{k}}g_{k+1}a^{k+1}_{i_{k}}&u_{n+1}v_{n+1}-f_{k}\\ \end{array}\right)$
$=$ $-\prod_{j=1}^{k}v_{i_{j}}\prod_{1\leq s<t\leq k}(a_{i_{t}}-a_{i_{s}})g_{k-1}+$
$\prod_{j=1}^{k}v_{i_{j}}\left(-g_{k+1}\det\left(\begin{array}[]{ccc}1&\ldots&1\\ a_{i_{1}}&\ldots&a_{i_{k}}\\ \vdots&\ddots&\vdots\\ a^{k-2}_{i_{1}}&\ldots&a^{k-2}_{i_{k}}\\ a^{k+1}_{i_{1}}&\ldots&a^{k+1}_{i_{k}}\\ \end{array}\right)+(u_{n+1}v_{n+1}-f_{k})\det\left(\begin{array}[]{ccc}1&\ldots&1\\ a_{i_{1}}&\ldots&a_{i_{k}}\\ \vdots&\ddots&\vdots\\ a^{k-2}_{i_{1}}&\ldots&a^{k-2}_{i_{k}}\\ a^{k}_{i_{1}}&\ldots&a^{k}_{i_{k}}\\ \end{array}\right)\right)$
$=$ $\prod_{j=1}^{k}v_{i_{j}}\prod_{1\leq s<t\leq k}(a_{i_{t}}-a_{i_{s}})\left(-g_{k-1}-g_{k+1}\det\left(\begin{array}[]{cc}\sigma_{1}({\mathcal{S}}_{k})&\sigma_{0}({\mathcal{S}}_{k})\\ \sigma_{2}({\mathcal{S}}_{k})&\sigma_{1}({\mathcal{S}}_{k})\end{array}\right)+(u_{n+1}v_{n+1}-f_{k})\sigma_{1}({\mathcal{S}}_{k})\right)$
$=$ $\prod_{j=1}^{k}v_{i_{j}}\prod_{1\leq s<t\leq k}(a_{i_{t}}-a_{i_{s}})\left(g_{k+1}\left(\sigma_{2}({\mathcal{S}}_{k})-\sigma_{1}({\mathcal{S}}_{k})^{2}\right)+(u_{n+1}v_{n+1}-f_{k})\sigma_{1}({\mathcal{S}}_{k})-g_{k-1}\right).$
Since $\prod_{j=1}^{k}v_{i_{j}}\prod_{1\leq s<t\leq k}(a_{i_{t}}-a_{i_{s}})\neq 0$ for any ${\mathcal{S}}_{k}=\{a_{i_{1}},\ldots,a_{i_{k}}\}\subseteq{\mathcal{S}}$, we deduce that $\det(M_{k+1})\neq 0$ if and only if
$g_{k-1}\neq g_{k+1}\left(\sigma_{2}({\mathcal{S}}_{k})-\sigma_{1}({\mathcal{S}}_{k})^{2}\right)+(u_{n+1}v_{n+1}-f_{k})\sigma_{1}({\mathcal{S}}_{k}).$
Case 2: $M_{k+1}$ does not contain the last column of the matrix given in (43). Hence, by an argument similar to Case 1, we get that
$\det(M_{k+1})=\prod_{j=1}^{k+1}v_{i_{j}}\prod_{1\leq s<t\leq k+1}(a_{i_{t}}-a_{i_{s}})\left(g_{k+1}\sigma_{2}({\mathcal{S}}_{k+1})-g_{k-1}\right)$
and $\det(M_{k+1})\neq 0$ if and only if
$g_{k-1}\neq g_{k+1}\sigma_{2}({\mathcal{S}}_{k+1}).$
Combining Cases 1 and 2, it follows from Theorem 17 and Lemma 10 that $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if
$g_{k-1}\notin\left\{g_{k+1}\left(\sigma_{2}({\mathcal{S}}_{k})-\sigma_{1}({\mathcal{S}}_{k})^{2}\right)+(u_{n+1}v_{n+1}-f_{k})\sigma_{1}({\mathcal{S}}_{k}):~{\mathcal{S}}_{k}\subseteq{\mathcal{S}}\right\}\cup\left\{g_{k+1}\sigma_{2}({\mathcal{S}}_{k+1}):~{\mathcal{S}}_{k+1}\subseteq{\mathcal{S}}\right\}.$
This completes the proof. ∎
Remark 2.
We can see that the condition $g_{k+1}\in\mathbb{F}_{q}^{*}$ in Theorem 20 is not necessary for the proof. In other words, if we fix $g_{k+1}=0$ and let
$\mathbf{g}=(g(a_{1}),\ldots,g(a_{n}),u_{n+1})\in\mathbb{F}_{q}^{n+1},\quad g(x)=g_{k-1}x^{k-1}+f(x)\in\mathbb{F}_{q}[x],$
then the proof of Theorem 20 shows that $(\mathbf{v},v_{n+1})\star\mathbf{g}$ is a deep hole of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ if and only if
$g_{k-1}\notin\{0\}\cup\left\{(u_{n+1}v_{n+1}-f_{k})\sigma_{1}({\mathcal{S}}_{k}):~{\mathcal{S}}_{k}\subseteq{\mathcal{S}}\right\}\Leftrightarrow\begin{array}[]{l}~~g_{k-1}\neq 0~{\rm and}~u_{n+1}v_{n+1}=f_{k},\\ {\rm or}~g_{k-1}\neq 0,~u_{n+1}v_{n+1}\neq f_{k},~{\rm and}\\ \quad~\sigma_{1}({\mathcal{S}}_{k})\neq g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1},~\forall~{\mathcal{S}}_{k}\subseteq{\mathcal{S}}.\end{array}$
Note also that the condition $\sigma_{1}({\mathcal{S}}_{k})\neq g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1}~{\rm for~any}~{\mathcal{S}}_{k}\subseteq{\mathcal{S}}$ is equivalent to saying that
${\mathcal{S}}~{\rm is~an}~(n,k,g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1}){\text{-set}}.$
Therefore, we recover exactly the same criterion as in Theorem 19.
However, we highlight that the proof of Theorem 19 reveals more connections between extended Han-Zhang codes and EGRS codes as well as Roth-Lempel codes, thereby contributing to the subsequent subsection.
IV-B More non-GRS MDS codes and their connections with Roth-Lempel codes
As indicated in Remark 2, the proof of Theorem 19 inspires us to explore more connections between extended Han-Zhang codes and EGRS codes as well as Roth-Lempel codes. By combining (11) and Lemma 9, we then expect to derive more non-GRS MDS codes and examine their relationships with Roth-Lempel codes.
First, under the conditions described in Theorem 17.1), we obtain a class of $n+2,k+1,n-k+2]_{q}$ non-GRS MDS codes that are monomially equivalent to certain Roth-Lempel codes.
Theorem 21.
Suppose that $\mathbf{g}=(g(a_{1}),\ldots,g(a_{n}),u_{n+1})\in\mathbb{F}_{q}^{n+1}$, where $g(x)=g_{k-1}x^{k-1}+f(x)\in\mathbb{F}_{q}[x]$ with $g_{k-1}\in\mathbb{F}_{q}^{*}$ and $f(x)\in\mathcal{V}_{k}$. Let $\mathcal{C}$$^{\prime}$ be the linear code generated by $\left(\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right),~\infty_{k+1}\right)$, where $G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}$ is a generator matrix of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ given in (10). If $\mathcal{S}$} is $k$-zero-sum free and $u_{n+1}v_{n+1}=f_{k}$ with $v_{n+1}\in\mathbb{F}_{q}$, then $\mathcal{C}$$^{\prime}$ is an $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code and it is monomially equivalent to the Roth-Lempel code ${\mathrm{RL}}_{k+1,0}({\mathcal{S}})$.
Proof.
From Lemmas 4 and 5 as well as Theorem 16, the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ is an $n+1,k,n-k+2]_{q}$ non-GRS MDS code with covering radius $\rho({\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty))=n-k+1$. It then follows from Lemma 9 and Theorem 17 that $\mathcal{C}$$^{\prime}$ is an $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code under the given conditions.
Note also that the matrix $\left(\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right),~\infty_{k+1}\right)$ is row equivalent to the matrix
$\left(\begin{array}[]{cccccc}v_{1}&\ldots&v_{n}&0&0\\ v_{1}a_{1}&\ldots&v_{n}a_{n}&0&0\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ v_{1}a_{1}^{k-2}&\ldots&v_{n}a_{n}^{k-2}&0&0\\ v_{1}a_{1}^{k-1}&\ldots&v_{n}a_{n}^{k-1}&g_{k-1}^{-1}(u_{n+1}v_{n+1}-f_{k})&g_{k-1}^{-1}\\ v_{1}a_{1}^{k}&\ldots&v_{n}a_{n}^{k}&1&0\end{array}\right),$ (50)
which implies that $\mathcal{C}$$^{\prime}$ is monomially equivalent to the linear code generated by the matrix
$\left(\begin{array}[]{cccccc}1&\ldots&1&0&0\\ a_{1}&\ldots&a_{n}&0&0\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ a_{1}^{k-2}&\ldots&a_{n}^{k-2}&0&0\\ a_{1}^{k-1}&\ldots&a_{n}^{k-1}&0&1\\ a_{1}^{k}&\ldots&a_{n}^{k}&1&0\end{array}\right)$
since $u_{n+1}v_{n+1}=f_{k}$. Combining with (9), we immediately conclude that $\mathcal{C}$$^{\prime}$ is monomially equivalent to the Roth-Lempel code ${\mathrm{RL}}_{k,0}({\mathcal{S}})$. This completes the proof. ∎
The following theorem yields another two classes of non-GRS MDS $n+2,k+1,n-k+2]_{q}$ codes with respect to the conditions described in Theorems 21.2) and 22.
Theorem 22.
Suppose that $\mathbf{g}=(g(a_{1}),\ldots,g(a_{n}),u_{n+1})\in\mathbb{F}_{q}^{n+1}$, where $g(x)=g_{k+1}x^{k+1}+g_{k-1}x^{k-1}+f(x)\in\mathbb{F}_{q}[x]$ with not both $g_{k+1}$ and $g_{k-1}$ equal to 0, and $f(x)\in\mathcal{V}_{k}$. Let $\mathcal{C}$$^{\prime}$ be the linear code generated by $\left(\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right),~\infty_{k+1}\right)$, where $G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}$ is a generator matrix of the extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ given in (10). Then the following statements hold.
  1. 1)
    If $g_{k+1}=0$, $g_{k-1}\neq 0$, $\mathcal{S}$} is both $k$-zero-sum free and an $(n,k,g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1})$-set, and $u_{n+1}v_{n+1}\neq f_{k}$ with $v_{n+1}\in\mathbb{F}_{q}$, then $\mathcal{C}$$^{\prime}$ is an $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code.
  2. 2)
    If $g_{k+1}\neq 0$, $\mathcal{S}$} is $k$-zero-sum free, and
    $g_{k-1}\notin\left\{g_{k+1}\left(\sigma_{2}({\mathcal{S}}_{k})-\sigma_{1}({\mathcal{S}}_{k})^{2}\right)+(u_{n+1}v_{n+1}-f_{k})\sigma_{1}({\mathcal{S}}_{k}):~{\mathcal{S}}_{k}\subseteq{\mathcal{S}}\right\}\cup\left\{g_{k+1}\sigma_{2}({\mathcal{S}}_{k+1}):~{\mathcal{S}}_{k+1}\subseteq{\mathcal{S}}\right\}$
    with $v_{n+1}\in\mathbb{F}_{q}$, then $\mathcal{C}$$^{\prime}$ is an $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code.
Proof.
By an argument similar to that in the proof of Theorem 21, we immediately obtain that $\mathcal{C}$$^{\prime}$ is an $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code by combining Theorems 21.2) and 22. This completes the proof. ∎
We have two remarks for Theorems 21 and 22 as follows.
Remark 3.
Unlike for Theorem 21, it is not straightforward to determine the monomial equivalence of non-GRS MDS codes $\mathcal{C}$$^{\prime}$ derived from Theorem 22 and Roth-Lempel codes. Taking Theorem 22.1) as an example and combining with (50), it is easily seen that the $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code $\mathcal{C}$$^{\prime}$ obtained is further monomially equivalent to the linear code generated by the matrix
$G_{{\mathcal{C}}^{\prime}}=\left(\begin{array}[]{cccccc}1&\ldots&1&0&0\\ a_{1}&\ldots&a_{n}&0&0\\ \vdots&\ddots&\vdots&\vdots&\vdots\\ a_{1}^{k-2}&\ldots&a_{n}^{k-2}&0&0\\ a_{1}^{k-1}&\ldots&a_{n}^{k-1}&1&1\\ a_{1}^{k}&\ldots&a_{n}^{k}&0&g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1}\end{array}\right).$ (57)
Consider ${\mathcal{S}}=\{1,2,5,6,9\}\subseteq\mathbb{F}_{11}$, ${{\mathbf{v}}}=(1,1,1,1,1)\in(\mathbb{F}_{11}^{*})^{5}$, and $g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1}=3$. Verified by Magma [5], we have the following facts:
  • $\mathcal{C}$$^{\prime}$ is a $7,4,4]_{11}$ non-GRS MDS code that is monomially equivalent to the Roth-Lempel code ${\mathrm{RL}}_{3,\delta}({\mathcal{S}})$ for $\delta\in\{3,7\}$;
  • $\mathcal{C}$$^{\prime}$ is a $7,4,4]_{11}$ non-GRS MDS code that is not monomially equivalent to any Roth-Lempel code ${\mathrm{RL}}_{3,\delta}({\mathcal{S}})$ for $\delta\in\{0,1,2,4,5,6,8,9,10\}$.
Remark 4.
Note that Theorem 21 produces a class of non-GRS MDS $n+2,k+1,n-k+2]_{q}$ codes that are monomially equivalent to Roth-Lempel codes ${\mathrm{RL}}_{k,0}({\mathcal{S}})$ for $\mathcal{S}$} being $k$-zero-sum free. Note also that Han and Fan showed in [13, Theorem 1] that ${\mathrm{RL}}_{k,0}({\mathcal{S}})$ is an NMDS code if and only if $\mathcal{S}$} contains a $k$-zero-sum subset. By combining Theorem 21 and [13], we immediately conclude that the class of Roth-Lempel codes ${\mathrm{RL}}_{k,0}({\mathcal{S}})$ is just a family of linear codes in which each code is either non-GRS MDS or NMDS. This fact also underscores the significance of the discussions presented in this subsection, as such codes hold particular interest and importance in coding theory, a point we have emphasized earlier.
In the following, we provide an algorithm for obtaining more non-GRS MDS codes based on Theorems 21 and 22.
Theorem 23.
More non-GRS MDS $n+2,k+1,n-k+2]_{q}$ codes can be obtained by Algorithm 2.
Proof.
The desired result follows immediately from Theorems 21 and 22. More details can be found in Algorithm 2. ∎
Input: An $n+1,k,n-k+2]_{q}$ extended Han-Zhang code ${\mathcal{C}}_{k}({\mathcal{S}},{{\mathbf{v}}},\infty)$
Output: An $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code $\mathcal{C}$$^{\prime}$
1 begin
2 for $g_{k+1},~g_{k-1}\in\mathbb{F}_{q}$ do
3    for $f(x)\in\mathcal{V}[x]$ do
4       for $u_{n+1},~v_{n+1}\in\mathbb{F}_{q}$ do
5          $g(x)\leftarrow g_{k+1}x^{k+1}+g_{k-1}x^{k-1}+f(x)$
6          ${\bf g}\leftarrow(g(a_{1}),g(a_{2}),\ldots,g(a_{n}),u_{n+1})$
7          $\mathcal{C}$$^{\prime}\leftarrow$ the linear code generated by the matrix $\left(\left(\begin{array}[]{c}G_{{\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)}\\ (\mathbf{v},v_{n+1})\star\mathbf{g}\end{array}\right),~\infty_{k+1}\right)$
8          $f_{k}\leftarrow$ the coefficient of $x^{k}$ in $f(x)$
9          if $g_{k+1}=0$, $g_{k-1}\neq 0$ and $u_{n+1}v_{n+1}=f_{k}$ then
10             return $\mathcal{C}$$^{\prime}$ is an $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code that is monomially equivalent to the Roth-Lempel code ${\mathrm{RL}}_{k+1,0}({\mathcal{S}})$
11          else if $g_{k+1}=0$, $g_{k-1}\neq 0$, $u_{n+1}v_{n+1}\neq f_{k}$ and $\mathcal{S}$} is an $(n,k,g_{k-1}(u_{n+1}v_{n+1}-f_{k})^{-1})$-set then
12             return $\mathcal{C}$$^{\prime}$ is an $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code
13          else if $g_{k+1}\neq 0$ then
14             $L_{1},~L_{2}\leftarrow$ two empty sets
15             for ${\mathcal{S}}_{k}\subseteq{\mathcal{S}}$ do
[…]
17
18               end for
19             for ${\mathcal{S}}_{k+1}\subseteq{\mathcal{S}}$ do
[…]
21
22               end for
23             if $g_{k-1}\notin L_{1}\cup L_{2}$ then
24                return $\mathcal{C}$$^{\prime}$ is an $n+2,k+1,n-k+2]_{q}$ non-GRS MDS code
25               end if
26
27            end if
28
29         end for
30
31      end for
32
33   end for
34
35 end
36
Algorithm 2 An algorithm for obtaining more non-GRS MDS codes from Theorems 21 and 22
We end this section with the following two examples.
Example 3.
Let $q=11$, ${\mathcal{S}}=\{3,4,5,6,7\}\subseteq\mathbb{F}_{11}^{5}$, and ${{\mathbf{v}}}=(1,1,1,1,1)\in(\mathbb{F}_{11}^{*})^{5}$. Note that $\mathcal{S}$} is $3$-zero-sum free. It then follows from Lemmas 4.1) and 5 that the extended Han-Zhang code ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ is a $6,3,4]_{11}$ non-GRS MDS code with generator matrix
$G_{3}=\left(\begin{array}[]{ccccccc}1&1&1&1&1&0\\ 3&4&5&6&7&0\\ 5&9&4&7&2&1\end{array}\right).$
From Theorem 16, we have $\rho({\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty))=4$. Furthermore, the following results hold.
  1. 1)
    Assume that $g_{4}=0$, $g_{2}=3$, and $f(x)=4x^{3}+10x+7\in\mathcal{V}_{3}$. Then $g(x)=4x^{3}+3x^{2}+10x+7$.
    • According to Theorem 19.1), we deduce that
      $(g(a_{1}),g(a_{2}),g(a_{3}),g(a_{4}),g(a_{5}),4)=(7,10,5,5,1,4)$
      is a deep hole of ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$.
    • Note also that $\mathcal{S}$} is a $(5,3,\delta)$-set for $\delta\in\{0,8,9,10\}$. It then follows from Theorem 19.2) that
      $(g(a_{1}),g(a_{2}),g(a_{3}),g(a_{4}),g(a_{5}),u_{6}v_{6})=\left\{\begin{array}[]{ll}(7,10,5,5,1,3),&{\rm if}~3(u_{6}v_{6}-4)^{-1}=8,\\ (7,10,5,5,1,8),&{\rm if}~3(u_{6}v_{6}-4)^{-1}=9,\\ (7,10,5,5,1,1),&{\rm if}~3(u_{6}v_{6}-4)^{-1}=10\end{array}\right$.
      are three deep holes of ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$. In addition, ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ does not have other deep holes related to the polynomial $g(x)$.
    • From Theorem 21, we directly obtain a $7,4,4]_{11}$ non-GRS MDS code generated by the matrix
      $\left(\begin{array}[]{cccccccc}1&1&1&1&1&0&0\\ 3&4&5&6&7&0&0\\ 5&9&4&7&2&1&0\\ 7&10&5&5&1&4&1\end{array}\right),$
      which is monomially equivalent to the Roth-Lempel code ${\mathrm{RL}}_{4,0}({\mathcal{S}})$.
    • From Theorem 22.1), we directly obtain three $7,4,4]_{11}$ non-GRS MDS codes ${\mathcal{C}}^{\prime}_{1}$, ${\mathcal{C}}^{\prime}_{2}$, and ${\mathcal{C}}^{\prime}_{3}$ generated by the matrices
      $\left(\begin{array}[]{cccccccc}1&1&1&1&1&0&0\\ 3&4&5&6&7&0&0\\ 5&9&4&7&2&1&0\\ 7&10&5&5&1&1&1\end{array}\right),~\left(\begin{array}[]{cccccccc}1&1&1&1&1&0&0\\ 3&4&5&6&7&0&0\\ 5&9&4&7&2&1&0\\ 7&10&5&5&1&3&1\end{array}\right),~{\rm and}~\left(\begin{array}[]{cccccccc}1&1&1&1&1&0&0\\ 3&4&5&6&7&0&0\\ 5&9&4&7&2&1&0\\ 7&10&5&5&1&8&1\end{array}\right),$
      respectively. Moreover, verified by Magma [5], we have the following facts:
      • ${\mathcal{C}}^{\prime}_{1}$ and ${\mathcal{C}}^{\prime}_{2}$ are not monomially equivalent to any Roth-Lempel code ${\mathrm{RL}}_{4,\delta}({\mathcal{S}})$ for any $\delta\in\mathbb{F}_{11}$;
      • ${\mathcal{C}}^{\prime}_{3}$ is monomially equivalent to the Roth-Lempel code ${\mathrm{RL}}_{4,\delta}({\mathcal{S}})$ for $\delta\in\{9,10\}$, but not monomially equivalent to other Roth-Lempel codes ${\mathrm{RL}}_{4,\delta}({\mathcal{S}})$ for $\delta\in\{0,1,2,3,4,5,6,7,8\}$.
  2. 2)
    Assume that $g_{4}=2$ and $f(x)=3x^{3}+5x+2$.
    • It can be checked that
[…]
      Hence, according to Theorem 20, for $u_{6}v_{6}=0$ and $g_{2}=8$, $i.e.,$ $g(x)=2x^{4}+3x^{3}+8x^{2}+5x+2$,
      $(g(a_{1}),g(a_{2}),g(a_{3}),g(a_{4}),g(a_{5}),u_{6}v_{6})=(2,7,4,7,1,0)$
      is a deep hole of ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ and ${\mathcal{C}}_{3}({\mathcal{S}},{{\mathbf{v}}},\infty)$ does not have other deep holes associated with $g(x)=2x^{4}+3x^{3}+g_{2}x^{2}+f_{1}x+f_{0}$ for any $g_{2},f_{0},f_{1}\in\mathbb{F}_{11}$.
    • From Theorem 22.2), we directly obtain a $7,4,4]_{11}$ non-GRS MDS code generated by the matrix
      $\left(\begin{array}[]{cccccccc}1&1&1&1&1&0&0\\ 3&4&5&6&7&0&0\\ 5&9&4&7&2&1&0\\ 2&7&4&7&1&0&1\end{array}\right).$
      Moreover, verified by Magma [5], this non-GRS MDS code is not monomially equivalent to the Roth-Lempel code ${\mathrm{RL}}_{4,\delta}({\mathcal{S}})$ for any $\delta\in\mathbb{F}_{11}$.
Example 4.
Let $q=13$, ${\mathcal{S}}=\{1,2,3,4,5,6,7,8,9,10,11,12\}\subseteq\mathbb{F}_{13}$, and ${{\mathbf{v}}}=(1,1,1,1,1,1,1,1,1,1,1,1)\in(\mathbb{F}_{13}^{*})^{12}$. Note that $\mathcal{S}$} contains a $5$-zero-sum subset, such as the $5$-subset $\{2,3,4,6,11\}$. It then follows from Lemma 4.2) that the extended Han-Zhang code ${\mathcal{C}}_{5}({\mathcal{S}},{{\mathbf{v}}},\infty)$ is a $13,5,8]_{13}$ NMDS code with generator matrix
$G_{5}=\left(\begin{array}[]{ccccccccccccc}1&0&0&0&0&10&7&9&6&8&2&0&3\\ 0&1&0&0&0&11&7&6&8&11&1&7&1\\ 0&0&1&0&0&12&5&12&2&12&6&7&5\\ 0&0&0&1&0&6&4&10&9&10&10&10&1\\ 0&0&0&0&1&1&4&3&2&12&8&3&3\end{array}\right).$
From Theorem 16, we have $\rho({\mathcal{C}}_{5}({\mathcal{S}},{{\mathbf{v}}},\infty))=9$. Furthermore, the following results hold.
  1. 1)
    Assume that $g_{6}=0$, $g_{4}=4$, and $f(x)=6x^{5}+3x^{3}+4x^{2}+2x+9\in\mathcal{V}_{5}$. Then $g(x)=6x^{5}+4x^{4}+3x^{3}+4x^{2}+2x+9$.
    • According to Theorem 19.1), we deduce that
      $(g(a_{1}),g(a_{2}),g(a_{3}),g(a_{4}),g(a_{5}),g(a_{6}),g(a_{7}),g(a_{8}),g(a_{9}),g(a_{10}),g(a_{11}),g(a_{12}),6)$
      $=$ $(2,10,3,5,8,3,11,10,5,7,12,6,6)$
      is a deep hole of ${\mathcal{C}}_{5}({\mathcal{S}},{{\mathbf{v}}},\infty)$.
    • Note also that $\mathcal{S}$} is not a $(12,5,4(u_{13}v_{13}-6)^{-1})$-set for any $u_{13},v_{13}\in\mathbb{F}_{13}$. It then follows from Theorem 19.1) that ${\mathcal{C}}_{5}({\mathcal{S}},{{\mathbf{v}}},\infty)$ has no other deep holes associated with the polynomial $g(x)$.
  2. 2)
    Assume that $g_{6}\neq 0$. Note that $\{\sigma_{2}({\mathcal{S}}_{6}):~{\mathcal{S}}_{6}\subseteq{\mathcal{S}}\}=\mathbb{F}_{13}$. For any $g_{6}\in\mathbb{F}_{13}^{*}$, it then follows that $\{g_{6}\sigma_{2}({\mathcal{S}}_{6}):~{\mathcal{S}}_{6}\subseteq{\mathcal{S}}\}=\mathbb{F}_{13}$, and hence, there is no $g_{4}\in\mathbb{F}_{13}$ such that
    $g_{4}\notin\left\{g_{6}\left(\sigma_{2}({\mathcal{S}}_{5})-\sigma_{1}({\mathcal{S}}_{5})^{2}\right)+(u_{13}v_{13}-f_{5})\sigma_{1}({\mathcal{S}}_{5}):~{\mathcal{S}}_{5}\subseteq{\mathcal{S}}\right\}\cup\left\{g_{6}\sigma_{2}({\mathcal{S}}_{6}):~{\mathcal{S}}_{6}\subseteq{\mathcal{S}}\right\}$
    for any $u_{13},v_{13},f_{5}\in\mathbb{F}_{13}$. According to Theorem 20, we immediately conclude that there is no deep hole related to the polynomial $g(x)=g_{6}x^{6}+g_{4}x^{4}+f(x)$ for any $g_{6}\in\mathbb{F}_{13}^{*}$, $g_{4}\in\mathbb{F}_{13}$, and $f(x)\in\mathcal{V}_{5}$.
In addition, since $\mathcal{S}$} contains a $5$-zero-sum subset, we cannot obtain more non-GRS MDS codes from Theorems 21 and 22.
V Conclusion
Extended Han-Zhang codes ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ constitute a class of linear codes where each code is either a non-GRS MDS code or an NMDS code, both of which have significant applications in communication, cryptography, and storage systems. Although various properties of these codes, including their dual codes, Schur squares, weight distributions, and (almost) self-dual properties, were investigated in [31], their decoding challenges have remained largely unexplored.
This paper aimed to bridge this gap, $i.e.,$ answering Problem 1. On the one hand, we fully examined the existence of $\ell$-ECPs $(\mathcal{A},\mathcal{B})$ for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ and provided explicit forms for $\mathcal{A}$ and $\mathcal{B}$, either via the subcode relationship or through a direct construction, as established in Theorems 13 and 14. Based on these $\ell$-ECPs, we proposed a decoding algorithm for ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ that can correct up to $\ell$ errors in $O(n^{3})$ time in Theorem 15 and Algorithm 1. On the other hand, we completely determined the covering radius of ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$ in Theorem 16 and characterized two classes of deep holes in Theorems 19 and 20, respectively. Building on these results, we constructed more non-GRS MDS codes with larger lengths and dimensions, and analyzed their monomial equivalence with Roth-Lempel codes in Theorems 21 and 22 as well as Remark 3. Furthermore, we provided an algorithm to generate these non-GRS MDS codes in Theorem 23 and Algorithm 2, and gave several examples to support our findings.
We highlight several research directions that warrant further exploration, summarized as follows.
  1. 1.
    In Theorem 15, we showed that the time complexity of Algorithm 1 is $O(n^{3})$. Similar to [22], [42], and [44], the extended Euclidean algorithm can also be applied to decode ${\mathcal{C}}_{k}({\mathcal{S}},{\bf v},\infty)$. As noted in [30], however, our decoding algorithm based on $\ell$-ECPs can achieve better performance than the extended Euclidean algorithm. Moreover, focusing on the system of equations
    $G_{\mathcal{B}}\cdot{\mathrm{diag}}({\bf y})\cdot G_{\mathcal{A}}^{T}\cdot{\bf s}^{T}={\bf 0},$
    which appears in Line 15 of Algorithm 1, we know from [10, Section \@slowromancapvi@.B] that there are two potential approaches to reduce the decoding complexity of Algorithm 1 to $O(n^{2})$ or even lower:
    • exploiting the possible regular structures in the matrix $G_{\mathcal{B}}\cdot{\mathrm{diag}}({\bf y})\cdot G_{\mathcal{A}}^{T}$;
    • reducing the size of the field that contains the entries of the matrix $G_{\mathcal{B}}\cdot{\mathrm{diag}}({\bf y})\cdot G_{\mathcal{A}}^{T}$.
    It would therefore be interesting to further investigate these two approaches as possible ways to reduce the decoding complexity of Algorithm 1.
  2. 2.
    As stated in Remark 3, the monomial equivalence between the $n+2,k+1,n-k+2]_{q}$ non-GRS MDS codes derived in Theorem 22 and Roth-Lempel codes remains unknown in the general case. Resolving this problem would directly contribute to the identification of new non-GRS MDS codes.